### v2

# Comprehensive Review of Deep Unfolding Techniques for Next-Generation Wireless Communication Systems

Abstract
The application of machine learning in wireless communications has been extensively explored, with deep unfolding emerging as a powerful model-based technique. Deep unfolding enhances interpretability by transforming complex iterative algorithms into structured layers of deep neural networks (DNNs). This approach seamlessly integrates domain knowledge with deep learning (DL), leveraging the strengths of both methods to simplify complex signal processing tasks in communication systems. To provide a solid foundation, we first present a brief overview of DL and deep unfolding. We then explore the applications of deep unfolding in key areas, including signal detection, channel estimation, beamforming design, decoding for error-correcting codes, sensing and communication, power allocation, and security. Each section focuses on a specific task, highlighting its significance in emerging 6G technologies and reviewing recent advancements in deep unfolding-based solutions. Finally, we discuss the challenges associated with developing deep unfolding techniques and propose potential improvements to enhance their applicability across diverse wireless communication scenarios.
Index Terms:
Deep learning (DL), deep unfolding, model-driven machine learning, wireless communication signal processing, ISAC, IRS, THz, mmWave.
I Introduction
The future mobile generations are expected to experience a surge in resource demand due to the exponential growth of connected devices. With the advent of 5G, billions of devices require seamless connectivity, intensifying competition for efficient spectrum allocation [1, 2, 3]. Each network has unique requirements and constraints, making resource allocation a complex challenge. Given the inherent limitations of available resources, problem formulation must prioritize flexibility, adaptability, and efficient utilization. Significant advancements in physical layer technologies have enhanced wireless communication systems, including massive multiple-input multiple-output (MIMO), intelligent reflecting surfaces (IRS), millimeter-wave (mmWave) communication, physical layer security, and non-orthogonal multiple access (NOMA). Additionally, orthogonal time-frequency space (OTFS) modulation is emerging as a promising alternative to orthogonal frequency-division multiplexing (OFDM), further improving spectral efficiency and resilience in next-generation networks.
In the communication system, traditional model-based analytical algorithms often employ iterative optimization techniques to achieve accurate solutions. Increasing the number of iterations improves performance as the algorithm converges closer to the optimal solution. Deriving an optimal solution in the form of a closed-form expression can be highly challenging in specific arbitrary communication scenarios. Hence, we resort to a sub-optimal solution in the traditional model-based method. Moreover, practical implementation on hardware imposes significant constraints on computational resources, latency, and power consumption, which limit the feasibility of deploying highly iterative methods. The classical iterative approach often relies on mathematically tractable models and algorithms for specific scenarios. Hence, adapting to dynamic and heterogeneous conditions, such as rapidly changing channels, interference, and diverse network configurations, can be challenging. Machine learning techniques offer a transformative approach by leveraging data-driven learning to extract patterns, adapt to new conditions, and approximate solutions efficiently. These methods can generalize across varying scenarios, learn optimal strategies from experience, and enable the design of intelligent communication systems that operate effectively in complex, real-world environments. These methods can potentially learn task-specific patterns and approximate solutions with significantly reduced computational complexity. Machine learning models can provide efficient and scalable alternatives to traditional iterative algorithms, enabling better performance under hardware limitations. This paradigm shift has opened new avenues for addressing complex problems in communication systems, bridging the gap between theoretical advancements and practical deployment.
Deep Learning (DL), a machine learning subclass maps the relationship between input and output without explicitly modeling it in the architecture [4]. It involves the use of deep neural networks (DNN) with multiple hidden layers (hence “deep”) to analyze and interpret data. Numerous architectures have been proposed such as multilayer perceptron (MLP), convolutional NN (CNN), deep reinforcement learning (DRL), generative adversarial networks (GANs), etc. DNNs have been applied to various components of wireless communication, including detection, channel estimation, beamforming designs, and power allocation [5, 6, 7, 8, 9]. DL architectures exhibit several inherent limitations. These include complex training procedures and a black-box nature. The training process for these models requires extensive datasets, considerable time, and intricate fine-tuning. The complexity of implementing DNN architectures often constrains their applicability in real-world systems. Therefore, DL models are highly sensitive to the specific configurations of wireless systems, such as the number of antennas, users, and operating frequencies, requiring retraining or reimplementation whenever these parameters change. These challenges are not typically encountered in conventional optimization techniques, underscoring the potential advantages of using traditional algorithms [10, 11].
Communication problems are inherently influenced by dynamic physical constraints that evolve in real-time. These constraints stem from factors such as the physical environment, hardware limitations, and the specific operational requirements of communication systems. Different communication scenarios impose varying demands on bandwidth, power consumption, and latency, which fluctuate with user density, data rates, and service types. Interpreting and addressing these constraints using conventional DNNs presents significant challenges. Traditional data-driven neural network architectures were developed primarily to process language and vision data, rendering them suboptimal for wireless communication tasks that involve channels, signals, and outputs governed by complex constraints. Although certain constraints can be handled through the application of specific activation functions, for example, employing the tanh function to enforce a constant modulus constraint in analog beamforming, standard activation functions generally lack the versatility to manage the intricate constraints characteristic of wireless communication systems. To mitigate these challenges, model-driven DL approaches facilitate the development of problem-specific network architectures and customized activation functions, optimized for the unique signal processing requirements of wireless communication applications [12].
Compared to the conventional fully-connected DL architectures, model-driven DL have fewer trainable parameters, implying less training time and computational complexity. Deep unfolding is an approach that falls into the category of model-driven DL [13]. The main motive behind developing a hybrid model (model-driven DL) for decision mapping, is to utilize the benefits of both optimization and data-driven DL. Model-driven DL is categorized into three frameworks as given in [13] : (1) learned optimizers [14, 15] (2) deep unfolding [16] (3) DNN-aided optimizers [17]. Here, we will focus on the second category i.e., deep unfolding. A deep unfolded network can be trained in an easier and faster fashion with smaller datasets than conventional DNNs. This is because deep unfolding incorporates domain knowledge from iterative algorithms into layers of DNNs, offering improved performance and interpretability [18]. The main idea in solving communication signal processing tasks lies in optimizing key hyperparameters of conventional algorithms based on data training. This includes elements such as step sizes in gradient-based methods, initial solutions, searching directions, criteria for early termination, and other relevant parameters. Deep unfolding uses suitable loss functions to train learnable parameters that are interpretable. Each iteration of the iterative algorithm is incorporated as a hidden layer of the DNN. These layers are then stacked to form a network, mimicking the optimization flow of algorithms. Furthermore, in many cases, the trainable parameters are independent of the system configurations, implying that the trained unfolding model can work for various scenarios.
I-A Related Works
The growing application of deep unfolding in communication systems highlights the need for a comprehensive survey in this area. Wang et al. in [19] presented a review of DL and deep unfolding approaches for addressing physical layer signal processing challenges, including MIMO detection, channel estimation, and channel decoding. The authors of [20] provided an overview of deep unfolding applications in transceiver design, MIMO detection, and CSI estimation and feedback. Similarly, in [21], the authors highlighted the role of DL in communication, with a focus on DNNs, GANs, RNNs, and end-to-end learning. This work also explored deep unfolding for Bayesian optimal estimation, using MIMO detection and channel estimation as case studies. The authors in [22] focused primarily on MIMO detection while exploring its use in multiuser MIMO precoding and belief propagation-based channel decoding. A survey on model-based, DL, and hybrid approaches in the context of wireless communication networks is provided in [23]. In [16], Monga et al. conducted a survey of deep unfolding algorithms specifically implemented for applications in signal and image processing. The authors in [24] reviewed the implementation of deep unfolding in signal processing tasks. A brief review on deep unfolding in MIMO detection is presented in [17]. The authors in [13] have presented a general framework for optimization that leverages the benefit of deep unfolding with signal detection as one of the case studies. In [12], a chapter is dedicated to model-based machine learning for communication. The deep unfolding networks have been discussed for signal detection, channel estimation, and channel coding. Machine learning-based optimization framework is discussed in [25] with an emphasis on deep unfolding in specific design paradigms such as signal detection. A comparative study between purely data-driven and deep unfolding is conducted with complete emphasis on hybrid beamforming optimization in [26]. An overview of existing surveys on deep unfolding is shown in TABLE I.
I-B Contributions
Existing surveys lack coverage of recent studies addressing critical 6G technologies, such as mmwave spectrum, new air interface that includes massive MIMO, IRS, next generation multiple access, and are often limited to specific application areas within the physical layer. Consequently, a comprehensive survey is needed to encompass the broad applications of deep unfolding techniques across various aspects of the physical layer in wireless communication. This paper provides an in-depth review of deep unfolding methods applied to the physical layer of wireless communication. The contributions are summarized below.
  1. 1.
    We provide a comprehensive overview of deep unfolding algorithms and their applications in physical layer wireless communication. Part A of the paper introduces the fundamentals of DL and deep unfolding to establish a foundation, outlining their general structures. We describe the DL architectures, including MLP, CNN, RNN, DRL, GNN, and GAN. Additionally, we detail a systematic, step-by-step approach for developing deep unfolding algorithms. Finally, we highlight the advantages and motivations for adopting hybrid approaches to tackle optimization tasks effectively.
  2. 2.
    We survey the application of deep unfolding in the areas of signal detection, channel estimation, precoder design, sensing and communication, decoding, power allocation, and physical layer security in Part B. Within each section, we present these areas with the key 6G technologies (like IRS, mmWave, etc.) used to aid readers to design frameworks that adapt to future technologies. Additionally, this survey explores the application of deep unfolding in integrated sensing and communication (ISAC).Also, we discuss the joint techniques in the physical layer in Part C. The different areas are joint detection and channel estimation, joint activity detection and channel estimation, joint channel estimation and precoder design, and joint detection and decoding.
  3. 3.
    We outline a set of challenges in developing deep unfolding models and propose future directions that could guide advancements in deep unfolding techniques to enhance their performance.
TABLE I: Existing Survey Papers on Deep Unfolding for Communication.
SNo. Reference Year Detection Channel Estimation Precoder Design Sensing and Communication Decoding in Error Correcting Codes Power Allocation Security
1 Deep Learning for Wireless Physical Layer: Opportunities and Challenges [19] 2017 $\checkmark$ $\checkmark$ $\checkmark$
2 Model-Driven Deep Learning for Physical Layer Communications [20] 2019 $\checkmark$ $\checkmark$
3 Deep Learning in Physical Layer Communications [21] 2019 $\checkmark$ $\checkmark$
4 Deep Unfolding for Communications Systems: A Survey and Some New Directions [22] 2019 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
5 Wireless Networks Design in the Era of Deep Learning: Model-Based, AI-Based, or Both? [23] 2019 $\checkmark$ $\checkmark$
6 Algorithm Unrolling: Interpretable, Efficient Deep Learning for Signal and Image Processing [16] 2021 $\checkmark$ $\checkmark$
7 Intelligent Radio Signal Processing: A Survey [27] 2021 $\checkmark$ $\checkmark$ $\checkmark$
8 Redefining Wireless Communication for 6G: Signal Processing Meets Deep Learning With Deep Unfolding [24] 2021 $\checkmark$ $\checkmark$ $\checkmark$
9 Model-Based Deep Learning: Key Approaches and Design Guidelines [17] 2021 $\checkmark$
10 Model-Based Deep Learning: On the Intersection of Deep Learning and Optimization [13] 2022 $\checkmark$
11 Machine Learning and Wireless Communications [12] 2022 $\checkmark$ $\checkmark$ $\checkmark$
12 Machine Learning for Large-Scale Optimization in 6G Wireless Networks [25] 2023 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
13 Artificial Intelligence-Empowered Hybrid Multiple-Input/Multiple-Output Beamforming: Learning to Optimize for High-Throughput Scalable MIMO [26] 2024 $\checkmark$
14 This Survey $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
I-C Paper Outline
The skeleton of this paper is depicted in Fig.  1.
Figure 1: Outline of the paper.
This survey paper is organized as follows:
  • Part A: We discuss the general structure of DNN and deep unfolding models. Section II covers the architecture of DNN and outlines the steps in designing a DNN. Section III presents the deep unfolding framework, which combines classical iterative algorithms with DNNs.
  • Part B: The applications of deep unfolding in different physical layer signal processing tasks are thoroughly reviewed. Specifically, Section IV explains signal detection with perfect knowledge of channel gains in the receiver, followed by channel estimation in Section V. Section VI covers the precoder design. Section VII reviews different deep unfolding approaches for ISAC. The decoding for error correction codes is discussed in Section VIII, followed by power allocation in Section IX and security aspects in Section X.
  • Part C : Section XI surveys the joint techniques in the physical layer. In this section, we discuss the application of deep unfolding in different joint techniques such as joint detection and channel estimation, joint activity detection and channel estimation, joint channel estimation and precoder design, and joint detection and decoding.
  • Part D : Section XII discusses the challenges in adopting the deep unfolding approach in wireless communication and the future scopes of research. Section XIII concludes the paper.
TABLE II: Acronyms and Corresponding Full Names.
Abbreviation Description Abbreviation Description
AMP Approximate message passing ADMM Alternating direction method of multipliers
BP Belief propagation BS Base station
BER Bit error rate CIR Channel impulse response
CNN Convolutional neural network CSI Channel state information
DL Deep learning DNN Deep neural network
DRL Deep reinforcement learning EM Expectation maximization
EP Expectation propagation FDD Frequency division duplexing
FFT Fast Fourier transform FDMA Frequency division multiple access
GAN Generative adversarial network GNN Graph neural network
HetNets Heterogeneous networks IRS Intelligent reflecting surface
IoT Internet of things ICI Inter-carrier interference
ISI Inter-symbol interference ISAC Integrated sensing and communication
ISTA Iterative shrinkage-thresholding algorithm JCEHB Joint channel estimation hybrid beamforming
JCDD Joint channel estimation detection and decoding LLR Log-likelihood ratio
LDPC Low-density parity check LPBN Long term passive beamforming
LMMSE Linear minimum mean square error LR Low resolution
LS Least squares LMS Least mean squares
LSTM Long short term memory MAP Maximum a posteriori
MDP Markov decision process MF Matched filtering
MO Manifold optimization MM Majorization minimization
MS Mobile station MMSE Minimum mean square error
MIMO Multiple input multiple output ML Maximum likelihood
MPD Message passing detector MSE Mean square error
MSER Minimum symbol error rate mMTC Massive machine type communication
NN Neural network NOMA Non-orthogonal multiple access
OAMP Orthogonal approximate message passing OFDM Orthogonal frequency division multiplexing
OMP Orthogonal matching pursuit OTFS Orthogonal time frequency space
PGA Projected gradient ascent PGD Projected gradient descent
RF Radio frequency RNN Recurrent neural network
RL Reinforcement learning RSMA Rate splitting multiple access
SSCA Stochastic successive convex approximation SABN Short term active beamforming network
SBRNN Sliding bidirectional recurrent neural network SCDMA Sparse code division multiple access
SE Spectral efficiency SGD Stochastic gradient descent
SIC Successive interfernce cancellation SISO Single-input single-output
SIMO Single-input multiple-output SINR Signal-to-interference and noise ratio
SNR Signal-to-noise ratio TAS Transmit antenna selection
TDD Time division duplexing THz Tera Hertz
UAV Unmanned aerial vehicle URLLC Ultra reliable low latency communication
WMMSE Weighted minimum mean square error ZF Zero forcing
I-D Notation and Acronyms
[…]
II Deep Neural Network: General Structure
Artificial intelligence originated with the famous Turing test in the 1950s, marking a breakthrough in machine thinking capability [28]. Over the decades, around the 1990s, machine learning formed a subclass under artificial intelligence. Machine learning requires manual feature extraction of structured data. Most machine learning algorithms can be categorized into supervised, unsupervised, and reinforcement learning [4].
Figure 2: Feedforward NN (MLP) architecture.
Supervised learning aims to learn how to map inputs to outputs to predict labels for new data. In unsupervised learning, the dataset includes many features that are not labeled. The algorithm tries to identify hidden patterns within the data without the guidance of labeled examples. The third category, reinforcement learning, consists of an agent that learns to make decisions based on the environment. Since there is no predefined dataset, the agent interacts with the environment and learns from the feedback it receives for every action [4],[29]. The output of a machine learning algorithm can be of two types: regression and classification.
In the 2010s, DL emerged as a crucial technology in designing intelligent networks with unstructured data [30]. Researchers have adopted DL techniques to solve complex problems. It has been widely used in computer vision, natural language processing, and image processing. This success has inspired researchers to use DL in the physical layer of wireless communication, especially in various signal-processing tasks.
An NN consists of interconnected layers of neurons that work together to learn patterns from the given data. Different types of artificial NN architectures have been developed for various learning tasks, such as feed-forward NNs, CNN, recurrent NN (RNN), graph NN (GNN), etc. Fig. 2 shows the feedforward NN architecture with $K$ ($K=2$ here) hidden layers and $N_{k}$ neurons in layer $k$, for all $k=0,\dots,K$. The architecture has an input layer, hidden layers, and an output layer.
We now discuss the steps in a feedforward NN:
  1. 1.
    The network begins with an input vector
    ${\bf{s}}=[s_{1},s_{2},\dots,s_{n}]$, at the input layer where $n$ is the number of input features.
  2. 2.
    For each neuron in a layer, the weighted sum of inputs is computed:
    $z^{(k)}_{j}=\sum_{i=1}^{n}w_{ji}^{(k)}s_{j}+b_{j}^{(k)},$
    where $w_{ji}^{(k)}$ is the weight connecting neuron $i$ to neuron $j$ in the $k^{\text{th}}$ layer and $b_{j}^{(k)}$ is the bias for neuron $j$ at the $k^{\text{th}}$ layer.
  3. 3.
    A non-linear activation function $\sigma(\cdot)$ is applied:
    $a_{j}^{(k)}=\sigma(z^{(k)}_{j}),$
    where $a_{j}^{(k)}$ is the activation output of the neuron $j$ in the layer $k$. Common activation functions include ReLU, sigmoid, tanh, softmax etc. [4].
  4. 4.
    At the output layer, after applying the weighted sum and activation process, the output is given by:
    $\hat{\bf{y}}=\sigma(\mathbf{W}^{K}\mathbf{s}^{K}+\mathbf{b}^{K})$
    where ${\bf{W}}$ is the weight matrix at the output layer and $\bf{b}$ is the bias vector at the output layer. The activation function, ($\sigma(\cdot)$) for the output often depends on the task, whether it is a classification or regression task.
  5. 5.
    The loss is computed to measure the network’s performance:
    $\mathcal{L}(\hat{\bf{y}},\bf{y})$
    where $\hat{\bf{y}}$ is the predicted output, and $\bf{y}$ is the true output. The common loss functions are mean squared error (MSE), binary cross entropy, categorical cross entropy, etc.
To train the NN, backpropagation is used by adjusting the parameters of the network, i.e., the weight and biases to minimize the loss. The parameters are updated using an optimization algorithm like stochastic gradient descent (SGD).
On the other hand, CNN is designed to process grid-like data by leveraging spatial structure. It uses layers of convolutional filters to extract features like edges followed by pooling layers to reduce spatial dimensions and retain key information [31]. The extracted features are passed through fully connected layers. One of the early applications of CNN was modulation classification based on signal constellation diagrams; the authors have used the CNN-based DL method [32]. They have developed and analyzed methods to represent modulated signals in data formats with grid-like topologies to conform with the CNN architecture.
RNN is another type of NN designed explicitly for processing sequential data. In RNN, the output from the previous step is fed as an input to the current step, incorporating sequential dependencies between the nodes. A recurrent neuron performs this function by maintaining a hidden state. Long short-term memory (LSTM) is one version of RNN [4]. The LSTM network was adopted into the NOMA framework in [33] to learn the channel characteristics of the NOMA system.
DRL integrates RL with DL to develop intelligent agents capable of making sequential decisions in complex environments. In DRL, an agent interacts with its environment to maximize cumulative rewards, learning optimal policies through exploration and exploitation. DNNs are employed to approximate value functions or policies, allowing DRL to efficiently handle high-dimensional state and action spaces. Key techniques such as Deep Q-Networks (DQN), policy gradient methods, and actor-critic frameworks enable agents to improve their performance iteratively through trial and error, guided by reward signals [4, 34]. In the domain of communications and networking, DRL has emerged as a powerful tool for addressing challenges such as resource allocation, spectrum management, and network optimization. Modern network architectures, including the Internet of Things (IoT), heterogeneous networks (HetNets), and unmanned aerial vehicle (UAV) networks, exhibit decentralized, dynamic, and autonomous characteristics, making them well-suited for reinforcement learning-based solutions [35].
The concept of using autoencoders to design end-to-end communication systems has been widely studied [36]. End-to-end learning trains a single model to map raw input directly to the desired output. It avoids intermediate steps or manual feature engineering. Autoencoders are neural networks that optimize both the transmitter and receiver together in this framework. In communication systems, the goal is to design the transmitter and receiver jointly for a given channel model. This helps achieve better performance, such as faster convergence and lower bit error rates [37].
GNNs are specialized networks designed for graph-structured data, where nodes are connected by edges [38]. Unlike traditional neural networks, GNNs work on non-Euclidean data and capture relationships between graph nodes. They learn rich representations of nodes, edges, or the whole graph by aggregating and transforming features from neighboring nodes. This ability to use graph structure helps GNNs perform better in tasks that rely on relational data.
In 2020, Goodfellow et al. proposed a new AI algorithm for generative modeling [39]. GANs use two neural networks: a generator and a discriminator. The generator creates realistic outputs, while the discriminator identifies if they are real or fake. By competing and learning together, GANs can produce high-quality outputs that are often indistinguishable from real data.
For further reading on application of deep learning in physical layer of communication, we invite readers to refer [40, 41, 42, 27].
III Deep Unfolding: General Structure
Deep unfolding, also known as algorithm unrolling [16], is a typical model-driven DL method [10],[20],[43, 44, 45, 46, 47, 48]. It has been frequently considered for various tasks such as signal detection, design of beamforming matrices, and many more optimization tasks in wireless communication. Owing to reduced training time and dataset, deep unfolding is one promising alternate approach to purely model-driven or data-driven.
Figure 3: Schematics of data-driven, classical and deep-unfolding approaches.
Deep unfolding integrates model-based approaches and domain knowledge into the design of DNN architectures. The training is carried out in an end-to-end manner. In deep unfolding, an iterative algorithm is mapped to DNN layers. Then, DL techniques such as SGD are used to learn the parameters. Fig. 3 shows a schematic diagram comparing the classical approach, data-driven approach, and the deep unfolding approach. Furthermore, Fig. 4 depicts the deep unfolding approach in greater details.
Figure 4: General framework of deep unfolding.
For any iterative algorithm, the update equation for each iteration is given as:
${\bf{s}}^{(t+1)}=h({\bf{s}}^{t};{\bf{\Psi}}^{t})$
where ${\bf{s}}^{t}\in{\mathbb{R}}^{n}$ denotes the variable vector to be updated in each iteration in the range $t=1,\dots,T$. The iterative function for the iterative algorithm is represented by $h(\cdot;\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. The trainable parameters, ${\bf{\Psi}}^{t}\in\mathbb{R}^{n}$ include the model parameters and the step size for the regularization coefficients of the algorithm [25]. Each iteration function $h(\cdot;\cdot)$ is mapped to a single network layer, and a finite number of layers are stacked together as shown in Fig. 4. To learn the trainable parameters ${\bf{\Psi}}^{t}$ in an end-to-end approach, the following optimization is carried out:
$\text{minimize}_{{\bf{\Psi}}_{T}}\mathcal{L}({\bf{s}}^{T}({\bf{\Psi}}_{T}))$
where $\mathcal{L}(\cdot)$ denotes the loss function for training, ${\bf{\Psi}}_{T}=\{\Psi^{t}\}_{t=0}^{T}$ is the set of trainable parameters of the total $T$-layer network and ${\bf{s}}^{T}(\cdot)$ is the output of the unfolded network. Feeding the data forward through the unfolded network with trainable parameters in the testing stage is similar to executing the parameter-optimized iterative algorithm for a finite number of iterations. Thus, the number of layers in deep unfolding models is fixed, resulting in a predefined computational complexity. Within these fixed layers, the model is trained to optimize the trainable parameters, enabling the problem to be addressed using a structured algorithm with parameters tailored for optimal performance.
The procedure to solve an optimization problem using a deep unfolding approach is outlined below.
  1. 1.
    Identify a principle algorithm that serves as the foundation for deep unfolding.
  2. 2.
    Select trainable parameters or specific blocks in the principle algorithm. Replace these with data-trained operations optimized through deep unfolding.
  3. 3.
    Define the loss function and represent the algorithm as a DNN. Each layer corresponds to an iteration of the original algorithm. Train the network using supervised or unsupervised methods. Since the number of iterations is predetermined, the network’s layers match these iterations. The number of layers can also be fine-tuned to balance performance and computational complexity.
TABLE III: Key technologies envisioned for 6G using deep unfolding.
Key technology Task Reference
Massive MIMO Detection [49],[50], [51], [52], [53], [44], [54], [55], [56], [57], [58], [59], [60], [61]
Channel estimation [62], [63], [64], [65], [66], [67]
Precoder Design [68], [69], [70], [71], [72], [73], [74], [75], [76], [77], [78], [79], [80], [81]
ISAC [82], [83], [84]
mmWave Channel Estimation [85]
Precoder design [86], [87], [88]
ISAC [89]
Modern Multicarrier Modulation OFDM Detection [90],[91], [92]
OTFS Detection [46]
OFDM Channel Estimation [93], [94], [95], [45], [96], [97], [98], [99], [47] [100]
OTFS Channel Estimation [101]
OFDM-ISAC [102], [103]
OTFS-ISAC [104]
Next Generation Multiple Access Detection [105], [106]
Precoder design [107]
Orbital Angular Momentum Detection [108]
IRS Channel Estimation [109], [110], [111], [112]
Precoder design [113], [114], [115], [116]
ISAC [117]
THz Precoder design [118], [119], [120]
Principle algorithms are typically iterative and based on optimization techniques. However, they are often complex, computationally demanding, and slow to converge. Deep unfolding improves these algorithms by leveraging data training. The key contributions of deep unfolding are:
  1. 1.
    Optimizing hyperparameters to enhance convergence, leading to better objective values or faster processing [26, 76, 84, 11].
  2. 2.
    Simplifying complex tasks, such as matrix inversion or decomposition, with lower computational complexity [121, 70, 122].
The first application is common in unfolded projected gradient ascent/descent (PGA/PGD) methods, while the second is widely used in unfolded weighted minimum mean square error (WMMSE) methods. Another approach involves learning the output at each layer, which corresponds to the iterative solution of the principal algorithm. This technique is widely applied in detection networks [123, 118, 49], following the PGA/PGD framework.
The main advantage of deep unfolding is its ability to incorporate domain knowledge into the network architecture. This improves performance while reducing complexity by using fewer parameters and a fixed number of layers. As a result, deep unfolding requires a smaller training dataset compared to generic DNNs. Next, we provide a comprehensive survey of deep unfolding algorithms designed for various signal processing tasks in wireless communication. Additionally, TABLE III lists key 6G technologies relevant to deep unfolding-based signal processing.
IV Signal Detection
This section discusses the application of deep unfolding techniques for signal detection. We segregate these signal detection techniques into various key technologies, as described in the following.
IV-A Massive MIMO
MIMO transmission uses multiple antennas at both the transmitter and receiver. It improves spectral and energy efficiency by utilizing a larger spatial domain. Adding more antennas (tens or hundreds) introduces the concept of massive MIMO [3]. However, detecting multiple symbols becomes more complex due to large-scale inter-symbol interference. Designing a low-complexity, high-performance detector for massive MIMO remains a challenge. For a MIMO communication system with $n_{\text{t}}$ transmit and $n_{\text{r}}$ receive antennas, the received signal can be given as:
$\mathbf{y}=\mathbf{Hx}+\mathbf{n}$ (1)
where $\mathbf{x}\in\mathbb{C}^{n_{\text{t}}}$ is the transmitted baseband signal vector, $\mathbf{H}\in\mathbb{C}^{n_{\text{r}}\times n_{\text{t}}}$ is the channel matrix and $\mathbf{y}\in\mathbb{C}^{n_{\text{r}}}$ is the received baseband signal vector. The detector extracts the desired transmitted information $\mathbf{x}$ from the received signal vector $\mathbf{y}$ with given $\mathbf{H}$ and the Gaussian noise vector $\mathbf{n}$.
The maximum a posteriori (MAP) and maximum likelihood (ML) estimations are Bayesian-based detection methods. These are the most commonly used classical detection techniques. The MAP rule applies when the transmitted signals have unequal a priori probabilities. The ML rule is used when these probabilities are equal [124]. ML minimizes the average error probability, providing an optimal solution. The ML detection rule is given as:
$\mathbf{\hat{x}_{\text{ML}}}=\text{arg}\min_{\mathbf{x}\in\mathcal{A}^{n_{t}}}||\mathbf{y}-\mathbf{Hx}||^{2}.$ (2)
This rule minimizes the Euclidean distance between $\bf{y}$ and $\bf{Hx}$. The minimization is performed over all possible combinations of the transmitted symbol vector $\bf{x}$ from the modulation alphabet $\mathcal{A}$. Since ML performs an exhaustive search and examines all possible signals, the computational complexity increases exponentially as $n_{\text{t}}$ increases.
MIMO detection can be linear or non-linear. Common linear detectors include matched filter (MF), zero-forcing (ZF), and minimum mean squared error (MMSE) [125]. These detectors require matrix inversion. Lattice reduction techniques can improve the performance of ZF and MMSE. Non-linear detectors, such as those using successive interference cancellation (SIC), perform interference estimation and removal in multiple stages [3], [126]. SIC and parallel interference cancellation are well-known techniques for this. In SIC, a detector like MF, ZF, or MMSE is used to detect the transmitted symbol in a data stream, leading to MF-SIC, ZF-SIC, or MMSE-SIC schemes. The data streams are arranged based on the received SNR. After detecting the symbol and estimating the channel coefficients, the interference caused by the detected symbol is estimated. This interference is then removed from the received signal, which is used to detect the next data stream in terms of received SNR. This process repeats until all symbols are detected. Sphere decoding is another non-linear detector [127]. It is an approximate ML detection method, where detection happens in a subspace of the modulation alphabet. The method searches within a bounded hypersphere around the received lattice point, using a radius for iterative symbol searching. The logarithmic likelihood ratio (LLR) is computed based on the symbol candidate found. This search region is smaller than the exhaustive ML search [128, 129].
The channel matrix for a MIMO fading channel can be modeled as a graph. The MIMO detection problem is framed as an inference task to calculate the posterior probability of system variables. Low-complexity algorithms are designed to compute these probabilities efficiently. The graph model combines probability and graph theory, showing the relationships between different nodes. Common graphical models include Bayesian belief networks, Markov random fields, and factor graphs [3]. Message-passing algorithms, like belief propagation (BP), are crucial for representing the MIMO system’s graph. Probability messages are exchanged between nodes representing transmitted symbols and received signals along the graph edges. The system is often modeled as a factor graph, where performance depends on the graph structure and algorithm convergence. Iterative algorithms, like message-passing detectors (MPD) with BP, approximate message passing (AMP), and channel hardening-exploiting message passing, provide solutions with lower complexity. Expectation propagation (EP) extends BP to handle larger constellations and higher-dimensional MIMO systems. EP approximates the posterior distribution of transmitted symbols iteratively, enabling efficient detection in polynomial time [130, 131].
The conventional detectors (linear and non-linear) involve computationally intensive matrix inversions or factorizations (like singular value decomposition or QR decomposition). Although these detectors yield nearly optimal solutions, they come at the cost of high complexity and high latency. Hence, with the aid of DL techniques, a deep unfolding approach has been adopted in literature to achieve low complexity and low latency.
PGD is an iterative optimization algorithm. Each step includes a gradient descent update to minimize a cost function, followed by a projection step to enforce constraints like sparsity or feasibility. In MIMO detection, PGD minimizes the error between the predicted and actual received signals while projecting the solution onto a feasible set, such as valid constellation points [132]. To achieve ML optimization, the PGD algorithm is unfolded into the DetNet architecture for MIMO detection [49, 133]. The gradient step size is learned across layers. The authors claimed that this approach reduces computational complexity compared to the sphere decoder and is robust to ill-conditioned channels. By leveraging the benefits of graphical models, He et al. proposed model-driven DL networks, OAMP-Net [10] and OAMP-Net2 [11], for MIMO detection with imperfect channel state information (CSI). OAMP-Net2 outperforms OAMP-Net by adapting to varying channel conditions. The orthogonal AMP (OAMP) algorithm is unfolded, and the LMMSE channel estimator is used to design the receiver. Learnable parameters, such as the step size for updating the prior mean and variance in the MMSE estimator, are trained using SGD. These parameters in OAMP-Net2 help compensate for channel estimation errors, improving performance compared to DetNet [49]. In [134], the authors introduced MMNet, a deep unfolded detector designed for a wide range of realistic channel matrices, including ill-conditioned and non-i.i.d. scenarios. MMNet is based on an iterative soft-thresholding algorithm, alternating between a linear detector and a non-linear denoising step. It balances architectural flexibility and computational complexity, outperforming DetNet [49] and OAMP-Net2 [11]. Overloaded massive MIMO systems, where the number of transmit antennas exceeds the number of receive antennas, require detectors with low complexity and good performance. In [61], the trainable projected gradient (TPG) algorithm was introduced for these systems. TPG uses a DL-aided iterative PGD method with trainable parameters, such as step size, optimized via backpropagation and SGD. TPG achieves lower computational cost and fewer trainable parameters compared to DetNet [49], OAMP-Net2 [11], and others. Additionally, a trainable detector based on the Hubbard–Stratonovich transformation is discussed in [135]. Takabe et al. avoided matrix inversion in this approach. Compared to the TPG detector [61] and OAMP-Net2 [10], the proposed network achieves faster training and lower execution cost.
In [136], the concrete MAP method is introduced, based on a continuous relaxation of the probability mass function for discrete random variables, as derived in [137]. This approach solves the MAP detection problem using continuous optimization instead of exhaustive search, approximating the MAP solution for any probabilistic nonlinear model. By unfolding the gradient descent algorithm, the detector provides soft outputs (detection probabilities), enabling better decoding compared to DetNet [49], OAMP-Net2 [10], and MMNet [134]. Yun et al. in [138] enhanced the unfolded PGD design by adding momentum components to the iterative hidden layers. Momentum helps guide parameter updates in the right direction and reduces oscillations during PGD optimization [139]. This modification improves convergence speed and overall performance, outperforming DetNet [49] and OAMP-Net2 [10, 11], while maintaining generality.
In wireless communication, physical layer signal processing involves continuous-time analog waveforms, which must be quantized for hardware implementation. Low-resolution one-bit analog-to-digital converters are commonly used in massive MIMO systems due to their high sampling rates, low cost, and low energy consumption [140]. To recover high-dimensional signals from noisy, low-resolution measurements, Khobahi et al. [141] developed a low-resolution detection network. This network extracts signals from noisy one-bit measurements without requiring prior knowledge of the channel matrix. It is based on unfolding the PGD method for symbol detection using the ML estimator, eliminating the need for CSI. The network outperforms traditional ML estimators [142] and DL networks like DeepSIC [143], even with small datasets of about 500 samples. In [59], a generalized PGD algorithm was proposed, eliminating the need for separate loss functions for projection and gradient steps. Instead of alternating after each step, projection is applied after several gradient steps, achieving a better balance. By unfolding this approach, the authors created a self-correcting auto detector that enhances performance [59].
Homotopy optimization transforms the original problem into a different one using a homotopy map. This method solves a series of simpler problems step by step, gradually progressing to the harder ones [144]. In MIMO detection, non-convex algorithms often struggle with the ML detection problem due to bad local minima [50]. By using a homotopy map to transition from easier to harder problems, detection performance is significantly improved. The study in [50] focuses on ML MIMO detection with one-bit quantized observations and binary symbol constellations. Using homotopy optimization, the authors apply a non-convex continuous reformulation of binary optimization [145] to address the detection problem. Their algorithm is designed to run efficiently and adapt to different channels. Building on this work, the authors introduced a proximal gradient-based homotopy method in [53]. This approach connects the structure of DetNet [49] with homotopy optimization for improved MIMO detection.
GNNs, as structured DNNs, excel in adaptability and flexibility. In MIMO detection, the graph structure of wireless networks makes GNNs particularly effective. Tan et al. [55] unfold MPD iterations to update messages at symbol nodes using prior and posterior LLR. Learnable parameters, such as prior probabilities and correction factors (damping, rescaling, offset), are optimized via mini-batch SGD to minimize the loss function. A greedy search determines the number of layers to balance performance and complexity. In [60], a low-complexity AMP-GNN model is proposed for massive MIMO detection. It outperforms OAMP-Net2 [10] by avoiding matrix inversion. For single-carrier FDMA systems, common in 5G to reduce peak-to-average power ratios, message-passing algorithms are computationally demanding. To address this, the unfolded network in [146] simplifies generalized AMP iterations [147], maintaining high performance even with channel mismatches. To enhance EP-based detection, [51] optimizes parameter selection using a modified moment-matching scheme, showing superior performance in high modulation cases compared to OAMP-Net2. In [52], approximate EP from [148] is unfolded, achieving better detection performance and robust convergence.
Liao et al. [44] incorporated interference cancellation into DNN layers using iterative embedding. Inspired by [149], their sequential detector trains residual error coefficients with MSE loss, reducing complexity by avoiding matrix inversion and relying on matrix addition and multiplication. Compared to DetNet [49, 133], which requires 90 layers for convergence, their model achieves convergence with just 8 layers, offering significant computational simplicity. Massive MIMO systems improve spectral efficiency (SE), coverage, and range but demand higher computational power. LcgNet, a model-driven DL architecture for massive MIMO detection, unfolds conjugate gradient descent iterations and learns step sizes through offline training [54]. Its quantized version, QLcgNet, minimizes memory usage with negligible performance degradation. LcgNet outperforms DetNet in terms of symbol error rate, computational complexity, and memory cost, making it a robust solution for massive MIMO systems.
The quantized version, QLcgNetV, reduces memory requirements while maintaining high detection performance. To enhance LcgNet, Olutayo et al. [150] incorporated a preconditioner during training, designing a preconditioned LcgNet. Preconditioning improves the condition number of a matrix [151], optimizing the receiver filter’s spectrum and minimizing eigenvalue spread. This enhancement significantly boosts the convergence speed compared to the original LcgNet [54]. Additionally, the dynamic conjugate gradient network utilizes self-supervised learning with forward error correction codes, adapting effectively to non-stationary, time-varying channels.
Sphere decoding achieves ML optimality but suffers from high computational complexity in massive MIMO systems. To address this, Nguyen et al. [123] proposed a fast DL-aided sphere decoding algorithm inspired by [152]. The DNN generates reliable initial candidates for sphere decoding, prioritizing candidates closer to the DNN’s output for testing. This approach accelerates sphere shrinkage, reducing computational cost. The sparse network architecture eliminates matrix inversion, relying only on element-wise matrix multiplications. The proposed method performs better in highly correlated channels than DetNet [49, 133], OAMP-Net2 [10, 11], MMNet [134], and LcgNet [54]. A MIMO detector based on annealed Langevin dynamics is introduced in [57]. It replaces the traditional two-step iterative framework (continuous-space optimization followed by projection) with a single-step approximation of the MAP estimator by sampling from Langevin dynamics. This approach shows superior performance in correlated channels compared to DetNet [49], OAMP-Net2 [11], and MMNet [134]. To mitigate the computational costs of matrix inversion in classical detectors like ZF and MMSE, Berra et al. [58] proposed unfolding accelerated Chebyshev successive over-relaxation and over-relaxation algorithms. This network requires fewer trainable parameters and features a fast, stable training scheme. The unfolded detector achieves better performance compared to DetNet [133], OAMP-Net2 [10], TPG-detector [61], MEPD [51], EP-Net [46], and other methods like [53, 44].
To avoid the computational burden incurred due to matrix inversion, the authors in [56] have proposed a sparse refinement architecture for symbol detection in uplink massive MIMO. The authors have used deep unfolding for various crucial components in matrix inversion, significantly reducing the detection complexity. For detection, the alternating direction method of multipliers (ADMM) algorithm is utilized in [153]. ADMM decomposes a complex optimization problem into simpler subproblems, which are solved iteratively. The algorithm combines the principles of dual ascent and the method of multipliers. Each iteration involves solving the subproblems for the primal variables in an alternating manner, followed by updating the dual variables (Lagrange multipliers) to enforce consistency with the original problem’s constraints [154]. This iterative process ensures that the subproblem solutions converge to the global solution of the main optimization problem. The authors in [155] developed an unfolded ADMM-based network for MIMO detection. They employed an inexact ADMM update [154, 156] to reduce complexity while achieving near-optimal performance. Compared to DetNet [49, 133], the proposed network requires fewer trainable parameters, resulting in improved performance in terms of symbol error.
IV-B Modern Multicarrier Modulation
An OFDM-based MIMO detector faces many challenges. The increased number of antennas at the transceiver exhibits enhanced spatial multiplexing but at the cost of increased inter-symbol interference (ISI). In standard OFDM multicarrier modulation, the cyclic prefix is inserted between OFDM blocks to mitigate ISI. Zhou et al. in [90], have unfolded a conjugate gradient method based OAMP detector for the OFDM-MIMO system. To address the issues faced in detecting coarsely quantized or low-resolution signals in OFDM-based MIMO systems, Shao et al. [91] have unfolded the expectation maximization (EM) algorithm. By unfolding the likelihood ascent search algorithm, Ullah et al. [92] are able to jointly detect and estimate the soft output in MIMO OFDM systems.
Li et al. [46] proposed a deep unfolding network for OTFS detection by converting the iterative EP algorithm into structured network layers. The core idea is to approximate the conditional probability distributions $p(\bf{y}|\bf{x},\bf{H})$ and $p(\bf{x})$ where $p(\cdot)$ represents the probability, and $\bf{y},\bf{x},\bf{H}$ denote the received signal vector, transmitted signal vector, and channel matrix, respectively. The trainable parameters in this network include the step size, damping factor, and offset values for the mean ($m$) and variance ($v$) of the probability distribution. During offline training, the number of layers in the unfolding network is selected to balance performance and computational complexity. To enhance convergence and prevent the network from being trapped in local optima, the Adam optimizer with warm restarts is employed. This technique adaptively adjusts the step sizes, improving both training stability and performance. In [101], an unfolded OAMP-based detector for OTFS is introduced to handle channel estimation errors. This detector is trained using imperfect CSI and optimizes three trainable parameters: step sizes for updating the linear estimate, nonlinear estimate, and noise variance.
IV-C Next Generation Multiple Access
A cellular network must simultaneously support a growing number of active subscribers within a limited time-frequency spectrum [157]. The current trend in wireless communication is to increase the signal bandwidth to achieve higher transmission rates. However, increasing bandwidth results in shorter symbol periods due to the inverse relationship between bandwidth and symbol period. Consequently, in a multi-path fading channel, the reduced symbol period exacerbates ISI. Multi-carrier modulation techniques like OFDM provide an effective solution to mitigate ISI by dividing a wideband signal into a set of orthogonal narrowband subcarriers [158]. This approach not only minimizes interference but also enhances robustness in multipath environments.
Multiple access techniques allow multiple users to share the same communication channel and resources, such as time, frequency, or code, to transmit data efficiently. Traditional multiple access schemes achieve this by dividing signaling dimensions orthogonally across time, frequency, or code [159]. This orthogonality enables low-complexity multi-user detection at the receiver. However, orthogonal multiple access cannot achieve the desired sum-rate capacity for a multiuser wireless system. To address this, NOMA has been introduced, allowing more users to be served than the number of orthogonal resource units by leveraging non-orthogonal resource sharing [157]. NOMA is categorized into power-domain NOMA and code-domain NOMA [160]. In power-domain NOMA, the transmitter superimposes information symbols into a single waveform using superposition coding. At the receiver, SIC is employed to iteratively decode signals until the desired signal is recovered.
A deep unfolded detector, named the learned preconditioned conjugate gradient descent network with SIC, is proposed for downlink MIMO-NOMA systems in [105]. To achieve a robust model with low dataset requirements, the algorithm incorporates a data transfer learning strategy. This approach adapts the model efficiently to variations in environmental factors, such as channel models, modulation schemes, or power allocation, ensuring consistent performance across diverse scenarios.
In sparse spread CDMA, each user is assigned spreading codes with only a few non-zero elements. This sparsity minimizes overlap and enables multiple users to transmit simultaneously on the same resources with minimal interference. In sparse CDMA, the spreading matrix contains sparse non-zero elements, as initially studied in [161]. Each user’s transmitted symbols are encoded using unique sparse signature sequences, which facilitate decoding at the BS [162]. Takabe et al. proposed a complex sparse trainable projected gradient detector (C-STPG) for SCDMA networks in [106]. The detector builds upon the PGD framework, delivering significantly improved detection performance compared to conventional LMMSE detectors. Simulation results demonstrate reduced computational complexity during execution and minimal training costs due to the limited number of trainable parameters. The authors also introduced gradual sparsification—a joint learning technique that designs sparse signature sequences using deep unfolding. This technique optimizes the weights of the signature matrix while reducing error during training through MSE loss. The trainable parameters include the step size for gradient descent ($\gamma$) and a softness parameter ($\alpha$) to adjust the strength of the soft projection.
IV-D Orbital Angular Momentum
Orbital angular momentum (OAM) MIMO multiplexing leverages electromagnetic waves with different orthogonal OAM modes [163]. These modes, theoretically infinite in number, form an orthogonal basis that can be used to multiplex distinct signals. A comprehensive review of OAM wave generation and detection is presented in [153]. In the context of OAM MIMO systems, Kamiya et al. [108] developed a signal detection algorithm specifically designed for systems employing uniform circular array antennas. These systems often face inter-mode interference, particularly when the transmitter and receiver antennas are misaligned, which degrades both MMSE capacity and detection performance. To address this issue, the ADMM algorithm is unfolded with parallel interference cancellation. This approach significantly enhances detection performance, even in scenarios where alignment imperfections introduce severe inter-mode interference.
In addition to detection, managing the peak-to-average power ratio (PAPR) presents a significant challenge in OAM multiplexing systems. To address this, Kamiya et al. [164] proposed a deliberate amplitude clipping method for OAM systems employing uniform circular array antennas. This approach involves applying amplitude clipping to digital baseband signals before OAM beamforming and pulse-shaping filtering. By doing so, it effectively reduces peak power, minimizes out-of-band radiation, and preserves the orthogonality of OAM modes. To counteract the bit error rate (BER) degradation introduced by clipping, the authors developed an iterative distortion recovery algorithm. Drawing inspiration from clipping noise cancellation techniques used in OFDM systems, the algorithm is implemented using deep unfolding. This implementation allows for trainable parameters that adaptively mitigate clipping-induced distortions, leading to enhanced detection performance and improved robustness in practical OAM systems.
V Channel Estimation
Estimating the channel in massive MIMO systems presents significant challenges, primarily because of the large number of antennas at the transceiver. The complexity of estimation grows with the scale of the system, and noise further exacerbates the issue, resulting in imperfect channel estimates and rendering estimation errors unavoidable [3], [165], [166], [167]. Pilot-based channel estimation techniques are commonly used to address these challenges. These techniques involve transmitting known pilot symbols, followed by processing at the receiver to estimate the channel’s characteristics and compensate for its effects. A block diagram illustrating the sequence of operations in a pilot-based channel estimation technique is provided in Fig. 5. It typically includes the generation and transmission of pilot symbols, reception at the receiver, estimation of the channel matrix using algorithms such as least squares (LS) or MMSE, and application of the estimated channel to process the received signals.
Figure 5: Block diagram for pilot-based channel estimation.
In the following, we discuss deep unfolding aided channel estimation methods.
V-A Massive MIMO
MMSE-based channel estimation is commonly used in MIMO systems to estimate the channel response in noisy environments. It minimizes the mean squared error between the actual channel and its estimate using the second-order statistics of the channel and noise. In massive MIMO systems with FDD, reducing CSI feedback overhead is challenging. A method in [62] combines deep learning and superposition coding to address this. Downlink CSI is encoded, superimposed on uplink user data, and sent to the base station. A multi-task neural network unfolds the MMSE criterion to reduce interference and reconstruct the CSI efficiently.
The iterative shrinkage-thresholding algorithm (ISTA) is commonly used for sparse channel estimation, enabling efficient CSI feedback. Each ISTA iteration includes a linear operation followed by a nonlinear soft-thresholding step, similar to the ReLU activation function [16]. In [63], Chen et al. proposed a trainable deep learning architecture for sparse recovery problems using an adaptive depth approach. Instead of a fixed number of layers, the method dynamically adjusts the depth based on the problem’s complexity. Highly sparse channels use fewer layers, while more complex channels require additional layers. This adaptive approach improves accuracy and lowers computational costs compared to traditional methods and fixed-depth networks. Results in [63] demonstrate that the method reduces reconstruction error and speeds up channel estimation, making it more efficient.
A scenario-adaptive CSI compression framework is proposed in [168], requiring minimal channel measurements for training. The authors introduced a CSI feedback architecture designed for efficient encoder updates in dynamic wireless environments, which is essential for downlink precoding in massive MIMO systems. The encoding network focuses on improving the robustness and efficiency of CSI compression. It employs a deep unfolding-based feedback network to iteratively refine the feedback process and enhance feature extraction. Similarly, the decoding network uses a deep unfolding structure inspired by the ISTA algorithm. This iterative approach improves CSI recovery accuracy, making the process more efficient and adaptable to changing conditions.
In massive MIMO systems, where the channel matrix is often sparse, the majorization-minimization (MM) approach is effective for addressing sparse channel estimation [169]. In [64], the authors address the critical challenge of downlink channel estimation in massive MIMO, essential for efficient communication. Traditional methods often rely on angular-domain channel sparsity to reduce pilot overhead but struggle in dynamic environments. To overcome this, the authors propose an online two-timescale joint uplink/downlink dictionary learning and channel estimation algorithm. This method leverages channel sparsity and angle reciprocity between uplink and downlink transmissions to adapt to dynamic scenarios and significantly reduce pilot overhead. The two-timescale strategy includes learning a long-term dictionary to capture spatial channel characteristics and frequently estimating short-term sparse channel vectors. The short-term channel estimation subproblems are solved using an unfolded MM algorithm, while the long-term dictionary optimization uses standard convex methods. Deep unfolding is employed to jointly optimize the dictionary learning and channel estimation process for both uplink and downlink. Simulations in [64] show that this approach reduces pilot overhead and outperforms existing methods in static and dynamic environments, especially in scenarios with mobile users.
The matching pursuit algorithm is a greedy iterative method often applied in sparse channel scenarios like massive MIMO. It identifies channel components (basis functions) by matching them to the residual, which is the difference between the observed signal and the signal generated by the current channel estimate. Yassine et al. [65] introduced a deep unfolded algorithm, called mpNet, based on the classical matching pursuit algorithm. This network estimates the multiuser massive MIMO channel in an unsupervised manner, initializing with a dictionary of imperfect steering vectors. The system operates in time division duplex (TDD) mode to leverage channel reciprocity. The network dynamically adjusts the number of layers based on incoming data, ensuring efficient performance in realistic environments.
The basis pursuit algorithm is a widely used method for sparse signal recovery, particularly in compressed sensing. It identifies the sparsest solution to an underdetermined system by minimizing the $\ell_{1}$-norm of the signal. This approach is crucial in scenarios where data is acquired under constraints, and the signal is assumed to be sparse in certain domains. In [66], the authors observed that existing sparse channel estimation schemes for downlink CSI often deliver suboptimal reconstruction due to the use of random measurement matrices. To address this, they proposed deep unfolding-based autoencoders that optimize the measurement matrix by unfolding the basis pursuit algorithm. This customized design enhances reconstruction performance in sparse channel estimation tasks.
Sparse Bayesian learning is another effective approach for channel recovery [170]. It estimates the non-zero elements of the channel matrix while utilizing prior knowledge of sparsity. Hu et al. [67] introduced a deep unfolding framework called deep deterministic policy gradient (DDPG) for multi-user MIMO systems. DDPG, a variant of deep Q-networks, is designed to handle problems in continuous action spaces [171]. Unlike traditional deep unfolding neural networks with fixed depth, the authors leveraged DDPG to dynamically adjust the network depth based on the complexity of the problem. This adaptive design enhances the efficiency and performance of sparse channel recovery in varying scenarios. The channel estimation problem is formulated using an off-grid sparse Bayesian learning model. The sparse Bayesian learning algorithm is unfolded layer by layer, introducing trainable parameters such as grid points and off-grid variables. Simulation results show that the proposed method significantly outperforms fixed-depth models, achieving a 30% reduction in the number of layers while improving normalized mean square error performance.
V-B mmWave Massive MIMO
mmWave massive MIMO operates in the high-frequency range of 30 GHz to 300 GHz, allowing higher dimensionality and the arrangement of many antennas in a compact space due to compressed wavelengths. The BS, equipped with numerous antennas, must acquire downlink CSI. However, the large amount of feedback required creates significant overhead, posing a challenge in mmWave massive MIMO systems. Compressed sensing techniques are widely used to reduce feedback overhead by exploiting channel sparsity [172]. The limited scattering properties of mmWave propagation and the power-focusing capability of lens antenna arrays transform the spatial domain channel into a sparse beamspace representation. Consequently, the channel estimation problem becomes a sparse signal recovery task, often addressed through compressed sensing methods. However, performing beamspace channel estimation in massive MIMO systems with limited RF chains remains a challenging task.
Compressive sensing aided deep unfolding is explored in [173], where the authors apply Onsager correction for AMP [174] and construct an unfolded model similar to [18]. The network in [175] uses deep unfolding to learn beamspace channel estimates. The authors demonstrate that this algorithm performs excellently with large antenna arrays and limited RF chains. Yi et al. in [176] introduced an unfolded network called LampResNet (Learned AMP with deep residual learning). In this model, the AMP algorithm is unfolded to estimate the channel, and deep residual learning is employed to mitigate channel noise. Normalized MSE analysis shows that the proposed networks outperform LAMP [173] and achieve performance similar to LDAMP [175]. However, exploiting channel sparsity does not always yield optimal estimation performance. To address this, the authors in [85] investigate mmWave channel sparsity in the beamspace. They formulate the estimation problem as a sparse signal recovery task and improve convergence speed by unfolding the iterative PGD algorithm. In this approach, the gradient descent step size is treated as a trainable parameter.
V-C Modern Multicarrier Modulation
For frequency-selective channels, OFDM is a preferred multicarrier modulation technique. It spreads the data across multiple orthogonal subcarriers in the frequency domain, helping to mitigate the effects of frequency-selective fading. The channel is typically modeled on a per-subcarrier basis.
To enable accurate CSI feedback by transforming the channel matrix into a sparse representation, Yiyun et al. [93] proposed a neural network consisting of compression and reconstruction modules for CSI compression and feedback. They designed the reconstruction module as an unfolded ISTA network and found that higher SNR improves reconstruction performance compared to the traditional BP method. Hu et al. [94] introduced an unfolded algorithm called the learnable optimization and regularization algorithm (LORA) to enhance CSI feedback speed and accuracy. In this approach, the conventional ISTA algorithm is unfolded, and the regularization term is learned through end-to-end training. They modeled the regularization term as a learnable, differentiable transform, further refined in [94]. By incorporating this learnable regularization module, the model better captures CSI characteristics and adapts to different levels of channel estimation error. To improve CSI feedback accuracy and efficiency in OFDM FDD-based massive MIMO systems, while reducing complexity at the user equipment, Yangyang et al. [95] proposed an unfolded ISTA network. Their study showed that the real and imaginary components of the channel differential term share similar correlations, simplifying network design. Unfolding ISTA significantly enhances the CSI encoder’s reconstruction performance.
In [45], Guo et al. proposed a two-stage low-rank CSI feedback scheme for a low-rank OFDM mmWave channel. In massive MIMO systems operating in FDD mode, CSI is estimated by unfolding the fast ISTA algorithm. Fast ISTA is an improved version of the conventional ISTA algorithm. A set of thresholds is derived from the input features of the network, and to reduce the uncertainty in CSI reconstruction, the iterative steps are mapped to layers of a neural network, with $L$ stages corresponding to $L$ iterations of fast ISTA. In [96], the authors explore quantized compressed CSI at the user equipment before feedback. They propose a deep unfolding-based bit-level CSI feedback method to mitigate the effects of accumulated quantization errors. Building upon ISTA, the authors use tied learned ISTA to reduce the number of learnable parameters, offering a faster and more accurate solution to the signal reduction problem. In tied learned ISTA, the trainable parameters are shared across all layers [177].
Jianqiao et al. [97] developed an unfolded inverse-free variational Bayesian learning method to reduce the high computational complexity of matrix inversion. The sparse channel recovery performance is enhanced by learning the a-priori parameters of the sparsely represented dictionary for channel estimation. In [98], based on approximate Bayesian inference and an extension of expectation consistent approximation, the authors proposed a generalized expectation consistent algorithm, an iterative message-passing approach for sparse channel estimation. In a lens-based beamspace massive MIMO system, this approach compensates for large channel attenuation in the mmWave band. The authors introduced a deep unfolded unsupervised learning network called the learned denoising-based generalized expectation consistent signal recovery network. This network unfolds the generalized expectation consistent signal recovery algorithm and formulates the wideband beamspace channel estimation problem as a compressed image recovery task. Stein’s unbiased risk estimator is used as the loss function. In [99], the authors leveraged angle-domain channel sparsity to reduce uplink pilot overhead during channel estimation. They designed an unfolded multiple-measurement-vectors learned AMP network by jointly training the phase shift network and channel estimator, integrating the channel sparsity for improved performance.
Underwater acoustic communication is widely used for long-range transmissions, reaching up to tens of kilometers [178, 179, 180]. The main challenges in underwater communication include frequency attenuation, time-varying multipath, and large Doppler effects. Zhao et al. [47] explored integrating the MMSE structure within the layers of a DNN to create an interpretable model. This approach aims to build an equalizer framework for improving underwater communication performance.
Figure 6: Analog, digital and hybrid beamformer [181].
Cognitive radio is a wireless technology that continuously monitors the spectrum to identify unused channels, or spectrum holes, allowing for more efficient spectrum usage. In [100], a wideband signal compressed spectrum reconstruction algorithm is proposed to detect spectrum holes across a large band using a sub-Nyquist sampling rate. The authors integrate ISTA into layers of a CNN, improving the performance in both the probability of detection and the probability of false alarm for spectrum hole detection.
V-D IRS
To mitigate the blockage effect and extend the connectivity range for mmWave communication, IRS is a key technology proposed to enhance both energy and SE. IRS consists of many reflecting elements that steer the propagation of electromagnetic waves by adjusting their reflection coefficients. However, to unlock the full potential of IRS, accurate channel estimation is required for both individual and cascaded links. This is a challenging task because IRS-based channel estimation involves estimating multiple channels simultaneously: the direct channels between the BS and each user, the channel between the BS and IRS elements, and the channel between IRS elements and users.
In [109], Jiguang et al. proposed a method for estimating the rank-deficient cascaded channel in an IRS-assisted SIMO system. They developed a deep unfolding network based on a gradient descent algorithm. The model structure is inspired by DetNet [49]. In massive MIMO, increasing the number of IRS elements and BS antennas enhances the system’s capacity. However, this leads to increased training overhead to obtain accurate channel estimates within the channel coherence time. The proposed network addresses this challenge by compensating for the computational complexity.
Parallel factor decomposition is an efficient technique for estimating multiple large channel matrices in massive MIMO communication systems. It allows a high-dimensional matrix to be decomposed into a linear combination of several rank-one matrices, making the estimation of low-rank matrices more computationally efficient [182]. In [110], the parallel factor decomposition is unfolded to estimate multiple large channel matrices in the uplink of an IRS-empowered multi-user MISO communication system.
Awais et al. [111] developed two networks based on the MAP estimation criterion. The first network is a deep denoising network designed for noise-dominated SNR regions, providing support at various noise levels. The second network is a deep unfolded half-quadratic splitting network that alternates between optimizing extrapolation and denoising for interference-dominated regions. As the SNR increases, the proposed network shows significant improvement in normalized MSE. Under severe channel variations, the network delivers a performance gain of 20% in the high SNR range.
Tsai et al. [112] proposed a two-stage learned AMP network with row compression to address the joint estimation problem of direct and cascaded channels in IRS-aided mmWave systems. By utilizing the deep unfolding technique, the network combines the advantages of compressive sensing and deep learning. This approach enables reliable channel estimation with reduced training, which is crucial for efficiently reconfiguring wireless environments under IRS reflection control.
VI Precoder Design
This section reviews the applications of deep unfolding for precoder design used along with key 6G technologies.
VI-A Massive MIMO
In MIMO systems, beamforming enhances signal transmission and reception by directing signals toward specific directions using multiple antennas. This improves the received SNR [181, 183]. There are three types of beamformers: analog, digital, and hybrid. Fig.  6 illustrates the architectures for analog, digital, and hybrid beamforming [181]. The hybrid beamformer system model is described in [184]. A digital beamformer assigns one RF chain per antenna. However, in large-scale arrays, it has high power consumption and complex architecture. Analog beamformers use time delay units or phase shifters for antenna weighting but struggle with inter-user interference in multi-user scenarios. To overcome these issues, hybrid beamforming was introduced [185, 186].
Beamforming design is an optimization problem where the objective is to maximize a network utility function, such as sum rate, mean rate, or harmonic rate. This optimization is subject to resource and hardware constraints, including power budget and constant modulus analog beamforming coefficients.
Consider a base station (BS) with $n_{\text{t}}$ antennas serving $M$ single-antenna users [27]. The received signal at user $m$ is given by:
$y_{m}=\boldsymbol{{h}}_{m}^{H}\sum_{m=1}^{M}\boldsymbol{{f}}_{m}x_{m}+n_{m}$ (3)
where, $x_{m}$ is the transmitted signal with an average power of $\mathbb{E}\{|x_{m}|^{2}\}=1$, $\mathbf{h}_{m}\in\mathbb{C}^{n_{t}\times 1}$ is the channel vector, $n_{m}\sim\mathcal{CN}(0,\sigma^{2})$ represents the AWGN noise at user $m$, and $\boldsymbol{f}_{m}\in\mathbb{C}^{n_{t}\times 1}$ is the beamforming vector for user $m$. The signal-to-interference-plus-noise ratio (SINR) at user $m$ is given by:
$\text{SINR}_{m}=\frac{\left|{\boldsymbol{{h}}_{m}^{H}\boldsymbol{{f}}_{m}}\right|^{2}}{\sum_{i=1,i\neq m}^{M}\left|{\boldsymbol{{h}}_{i}^{H}\boldsymbol{{f}}_{i}}\right|^{2}+\sigma^{2}}.$ (4)
This expression quantifies the received signal quality by measuring the desired signal power relative to interference and noise. The hybrid beamformer design problem can be formulated as the maximization of the sum rate, given by:
$(\mathcal{P}_{\text{HBF}})\quad\max_{\boldsymbol{F}^{\text{DB}},\boldsymbol{F}^{\text{AB}}}$ $\sum_{m=1}^{M}\log_{2}\left(1+\text{SINR}_{m}\right)$ (5)
s.t. $\left\lVert\boldsymbol{F}^{\text{AB}}\boldsymbol{F}^{\text{DB}}\right\rVert_{F}^{2}\leq P,\quad\boldsymbol{F}^{\text{AB}}\in\mathcal{F},$
$|\boldsymbol{F}^{\text{AB}}|^{2}=1.$
where $\boldsymbol{f}_{m}=\mathbf{F}^{\text{AB}}\boldsymbol{f}_{m}^{\text{DB}}$, $\mathbf{F}^{\text{AB}}\in\mathbb{C}^{n_{t}\times N_{\text{RF}}}$ is the analog precoding matrix with $N_{\text{RF}}$ RF chains, belonging to a codebook $\mathcal{F}$. The digital precoding vector for the $m^{\text{th}}$ user is $\boldsymbol{f}_{m}^{\text{DB}}\in\mathbb{C}^{N_{\text{RF}}\times 1}$, and the digital precoding matrix $\boldsymbol{F}^{\text{DB}}$ consists of column vectors $\boldsymbol{f}_{m}^{\text{DB}}$. These problems are nonconvex and NP-hard, meaning they lack closed-form solutions. The unit-modulus constraints on the analog precoding matrix make precoder design even more challenging. To address this, the algorithm unrolling technique combines model-driven and data-driven approaches. It optimizes beamforming matrices by integrating the interpretability of model-based methods with the adaptability of data-driven techniques. Shlezinger et al. [26] discuss the architectural constraints and challenges of hybrid beamforming optimization using both conventional and data-driven methods.
The classic iterative WMMSE algorithm is widely used for precoder design to maximize the system’s weighted sum rate. It assigns different weights to the MSE of each user based on priority or quality of service. By adjusting these weights iteratively, WMMSE balances user performance and improves overall efficiency. The algorithm provides closed-form, iterative update rules to solve a surrogate optimization problem with the same local optima as the sum-rate maximization problem [122]. Maximizing the weighted sum rate while satisfying a power constraint is a nonconvex problem [187]. The approach proposed in [187] exploits the separable structure of power constraints. To achieve a locally optimal beamforming strategy, the authors introduce a simple and efficient distributed inexact cyclic coordinate descent method.
A deep unfolded version of the iterative WMMSE algorithm is proposed for precoder design in multi-user MIMO systems [69]. The authors call this network architecture the iterative algorithm induced deep unfolding network (IAIDNN). To improve efficiency, WMMSE is reformulated to approximate matrix inversions. A generalized chain rule in matrix form helps visualize the gradient recurrence between adjacent layers during backpropagation. Using this rule, the gradients of trainable parameters are computed effectively. The WMMSE algorithm [122] involves matrix inversions, eigendecompositions, and bisection searches in its iterations. Since any MMSE solution requires matrix inversion, WMMSE also relies on it. To improve convergence while avoiding matrix inversion, a downlink beamforming method using an unfolded PGD-WMMSE algorithm for multi-user MISO systems is proposed in [68]. The authors replace matrix inversions with PGD and compute gradients in parallel. Unlike IAIDNN [69], this approach eliminates matrix inversion, making hardware implementation more practical. Additionally, it achieves a high weighted sum rate in fully loaded scenarios, where the number of users matches the number of transmit antennas at the BS.
In [69], learnable parameters are used to approximate matrix multiplication and inversion in WMMSE update rules. However, the network fails to exploit the inherent graph structure of wireless networks. This limitation is addressed in [70]. An unfolded WMMSE algorithm for transceiver design in multicell MU-MIMO interference channels with local CSI is proposed in [70]. The authors use graph convolutional networks (GCNs) to unfold WMMSE, enabling coordination in multicell MU-MIMO systems. Graph filters and convolutional networks approximate certain WMMSE variables using graph learners. However, the approach in [70] lacks flexibility in learning the exact transformation of WMMSE variables. To overcome this, [71] introduces a hybrid unfolded WMMSE algorithm for transmit beamforming in MU-MIMO ad-hoc networks. This method leverages GNNs to learn functional transformations of WMMSE parameters. Compared to the classical WMMSE algorithm, the proposed network reduces the number of layers, leading to faster convergence of the optimization problem. In [188], the joint optimization of precoders for multicell MU-MIMO systems is studied. The WMMSE algorithm is unfolded to enable distributed and synchronized precoding across cooperative BSs. Numerical analysis compares this approach with the networks in [69, 68]. The methods in [69, 68] rely on centralized optimization, leading to high complexity and backhaul overhead. In contrast, the proposed network in [188] reduces computational cost while maintaining efficiency.
The gradient descent algorithm is a widely used method for finding a stationary point. Shi et al. [72] designed a hybrid analog-digital transceiver using iterative gradient descent to obtain minimum symbol error rate (MSER) in massive MIMO systems. They unfolded the iterative gradient descent algorithm into a DNN, where learnable parameters serve as step sizes for the gradients. By using a fixed number of DNN layers with optimized step sizes, the approach reduces the number of iterations and speeds up convergence compared to conventional methods. The input to this MSER-based algorithm includes a user’s transmit beamforming matrix, an analog transmit beamforming matrix, an analog receive beamforming matrix, and a digital receive beamforming matrix. The step sizes guide the gradient descent update for each matrix. The gradients for these parameters are updated using SGD. The authors claim that their proposed network matches the performance of the MSER-based iterative gradient descent algorithm while significantly reducing complexity. Zhu et al. [73] proposed a beamforming learning architecture by unfolding a parallel gradient projection algorithm for MISO heterogeneous systems. Their design optimizes scenarios where network parameters, such as the number of antennas, BSs, or overall network size, differ from those used during training. Gradient-based methods have been successfully applied to beamforming optimization with reduced complexity. However, these approaches can be further improved using advanced optimization frameworks like PGA.
A fast and robust precoding network for multi-user MIMO systems with imperfect CSI and per-antenna power constraints is proposed in [74]. The objective function is formulated as a geometric program that minimizes the MSE between the transmitted and recovered symbol vectors at the user. This optimization leverages the duality between downlink and uplink MSE feasible regions. The authors map the iterative optimization process into a layer-wise deep unfolding framework, enabling fast and robust downlink precoding prediction.
PGA is a variant of the gradient ascent algorithm used to maximize an objective function under constraints. It updates the solution in the gradient direction to increase the objective function and then projects it back onto the feasible set. Lavi et al. [75] unfolded the PGA algorithm to optimize hybrid precoding for multiuser downlink MIMO, assuming perfect CSI availability. To handle noisy CSI, they implemented the projected conceptual mirror prox algorithm [189, 190]. The step sizes for updating the analog and digital precoder matrices are tuned using mini-batch SGD, a data-driven approach. The solution is computed within a fixed number of iterations, defined as network layers, ensuring a balance between interpretability and efficiency. In [76], Bilbao et al. proposed an efficient analog beamforming design to maximize the achievable rate of massive MIMO in-band full-duplex systems. The beamforming problem is non-convex due to constant modulus constraints on analog beamforming coefficients. To address this, they introduced an unfolded algorithm combining PGA with ZF projection onto convex sets, achieving faster convergence.
Wang et al. [191] proposed a universal algorithm for downlink massive MU-MIMO systems with finite-alphabet precoding to minimize inter-user interference. The algorithm is an improved version of the original ADMM framework. The authors observed that the standard ADMM struggles with nonlinear discrete optimization due to biased alternating updates. To address this, they introduced iterative discrete estimation [191], ensuring unbiased updates and better convergence. To reduce hardware costs and power consumption in multi-user massive MIMO, He et al. [77] developed an unfolded network using the IDE-based algorithm with finite-alphabet precoding. Additionally, constructive interference-based symbol-level precoding for multi-user MISO downlink transmission is discussed in [78]. In [79], the conjugate gradient algorithm is unfolded for constant envelope precoding to reduce computational overhead. The Riemannian manifold optimization (MO)-based conjugate gradient descent is implemented across network layers. Simulations show that this approach effectively suppresses multi-user interference, improving overall performance.
In [192], the authors use deep unfolding to optimize user scheduling, BS clustering, and beamforming in a dense C-RAN system. Reducing the number of BSs per user helps ease backhaul congestion. The goal is to maximize the weighted sum rate while minimizing BS clustering, all within power constraints. The PGD algorithm is unfolded with step sizes learned in an unsupervised manner. This approach avoids matrix inversion and bisection search, reducing iterations and computational complexity.
In [80], the interior point method is unfolded in an unsupervised manner for precoding design. It leverages interference in multi-user MISO systems to minimize power while meeting SINR constraints. In [81], the authors address the challenges of mutual coupling in beamforming. They propose mitigating its effects through covariance matrix reconstruction by unfolding an augmented coprime array.
VI-B mmWave
mmWave wireless communication, operating in the 30–300 GHz band, offers abundant spectral resources to alleviate the bandwidth crunch at sub-6 GHz frequencies [157]. However, mmWave signals suffer from severe propagation and penetration losses. To compensate for these losses, large-scale antenna arrays with tens to hundreds of elements are required to achieve high power gains. The short wavelength (1–10 mm) of mmWave signals allows compact integration of these arrays into small transceivers. These features introduce unique challenges in designing physical-layer transmission algorithms [193]. mmWave MIMO systems rely on beamforming to mitigate multi-user interference while enhancing diversity and spectrum efficiency.
Typically, the beam selection and digital precoding matrices are designed separately, leading to suboptimal solutions. To address this, Hu et al. [86] proposed a joint design approach combining DL and model-driven techniques. DRL optimizes beam selection, while a deep unfolding neural network, based on iterative WMMSE, optimizes the digital precoding matrix. The goal is to maximize the system sum rate while satisfying beam selection constraints. The DRL and deep unfolding networks are trained sequentially, with one network’s output serving as the input to the other. Backpropagation and SGD are used for training and parameter updates.
Lin et al. [87] studied a linear MMSE transceiver design for uplink massive MU-MIMO with one-bit ADCs. To minimize the system MSE by optimizing the precoder at each user equipment, the PGD algorithm is unfolded. The authors focused on training only the step size in each PGD iteration, significantly reducing the training overhead. In [88], the authors combined deep generative models with unfolding techniques to achieve near-optimal hybrid beamforming while minimizing feedback and computational complexity. They represented the channel as a low-dimensional manifold using a GAN, enabling efficient optimization of digital and analog precoders. The hybrid beamforming problem is formulated as a bilevel optimization, where an iterative rule updates both precoders. The upper-level optimization refines the digital precoder, leveraging full knowledge of the lower-level task constraints. Simultaneously, the lower-level optimization adjusts the analog precoder based on the digital precoder’s decisions. This structured optimization approach ensures efficient beamforming design with reduced computational costs.
VI-C Next Generation Multiple Access
Rate-Splitting Multiple Access (RSMA) is a flexible multiple access technology that merges the strengths of space division multiple access (SDMA) and NOMA. It uses linearly precoded rate-splitting at the transmitter and SIC at the receivers. RSMA balances interference management by partially decoding some interference while treating the rest as noise, bridging the gap between SDMA (which treats all interference as noise) and NOMA (which fully decodes interference) [194]. In [107], the authors applied fractional programming with hyperplane fixed-point iteration to design the RSMA beamformer. The algorithm is unfolded into a neural network, replacing its inner loop with a small-scale DNN to predict Lagrangian dual variables. Simulations show that the proposed approach achieves high performance while reducing computational complexity.
VI-D IRS
An IRS consists of a programmable array of passive elements that dynamically adjust the phase shifts of incoming signals. This enables signals to align in phase at the receiver, enhancing reception. IRS can mitigate interference, create alternative propagation paths to bypass obstacles, and extend signal coverage to previously unreachable areas. The phase shifts are optimized to reflect signals coherently toward the intended user while minimizing interference in other directions. Therefore, jointly optimizing IRS phase control and transmit beamforming enhances the performance of IRS-assisted systems [195].
With the aid of IRS, Chen et al. [113] addressed the joint design of the hybrid beamforming matrix and the phase-shifting matrix of the IRS elements. They proposed the deep unfolded WMMSE MO algorithm (DU-WMMSE-MO) using a deep unfolding methodology. The algorithm showed improved performance in IRS phase-shift design, utilizing two loops to jointly update the step size of the beamforming and IRS matrices. Compared to the Armijo backtracking line search method used in conventional WMMSE-MO [196], DU-WMMSE-MO provides more flexibility but with a higher runtime. Training is performed in an unsupervised manner, with a loss function that aligns with the objective function. The trainable parameters include the weights for the precoder, phase configuration of the IRS elements and combiner, and biases for the precoder and phase-shift matrix of the IRS.
An IRS-assisted multiuser MIMO full-duplex system is explored in [114]. The active beamforming matrices at the BS and uplink users, along with the passive beamforming matrix at the IRS, are jointly optimized to maximize the weighted sum-rate of the system. To address the high computational complexity of matrix inversion in the conventional stochastic successive convex approximation (SSCA)-based optimization algorithm, a deep-unfolding neural network is proposed. Fig. 7 shows two such netwroks from [114] . The network consists of a long-term passive beamforming network (LPBN) and a short-term active beamforming network (SABN). The learnable parameters of the LPBN and SABN are the IRS phase shift or passive beamforming vector $\boldsymbol{\theta}$ and the active beamforming vector $\boldsymbol{\Psi}$, respectively.
Figure 7: Two networks LPBN and SABN [114].
In [115], Ming-Min et al. introduced a two-timescale transmission approach to reduce the signal processing complexity and channel training overhead for the IRS-aided multiuser MISO system. To minimize the average transmit power at the AP while ensuring the individual quality of service constraints for the users, a primal-dual decomposition-based two-timescale hybrid beamforming method is proposed and unfolded. This approach allows for efficient beamforming optimization by separating the optimization into two timescales, improving both computational efficiency and performance.
In [197], a low-complexity IRS-assisted downlink MISO system is designed using an unfolded block coordinate descent algorithm to address the joint optimization problem of transmit beamforming and IRS phase shifts. The multi-ratio fractional programming structure of the objective function is handled by decoupling the problem into two subproblems: one for optimizing the beamforming at the base station and another for optimizing the phase shifts of the IRS. This is achieved through the use of Lagrange dual transformation, effectively simplifying the optimization process. The numerical analysis validated the achieved weighted sum rate is comparable to [198] but with reduced computational complexity. Furthermore, the authors in [199], have developed an unfolded network to address the joint beamforming task for an IRS-assisted MISO system. The WMMSE and power iteration algorithm are used as an alternating optimization strategy to derive the local optimal solution for the active and passive beamforming. The proposed network reduce the computational runtime to approximately 3% of the block coordinate descent algorithm in [198].
In [116], an IRS-assisted MISO covert symbiotic radio communication system is explored, and a deep unfolding algorithm based on gradient descent is proposed for beamforming design. The goal is to optimize the achievable transmission rate of the primary transmitter by refining both active and passive beamforming strategies while satisfying two key constraints: (1) ensuring covertness and (2) maintaining the SNR of the secondary transmission. The complex optimization problem, along with its constraints, is transformed into the Lagrangian dual domain. Each layer in the deep unfolded network consists of four modules that update: (1) the active beamforming, (2) the passive beamforming, and (3) two Lagrange multipliers corresponding to the two constraints.
Gangyong et al. in [200] have designed the IRS phase shift matrix by maximizing the sum path gain of the channel. This maximization is achieved by unfolding the Riemannian gradient descent algorithm. The network takes channel matrices and the initial reflection coefficients as inputs and outputs the optimized IRS coefficients. By configuring the IRS element phase shifts, the system effectively directs signal propagation paths to maximize the sum-path-gain, which is a key metric for evaluating beamforming quality and learning optimal IRS phase shifts to enhance the overall beamforming pattern.
VI-E THz
6G wireless networks are expected to achieve terabit-per-second (Tbps) data rates to support ultra-reliable low-latency communication. To enhance spectral and energy efficiency, THz-enabled massive MIMO has emerged as a promising solution, leveraging the abundant bandwidth at THz frequencies to significantly boost data rates. Additionally, hybrid beamforming offers a cost-effective approach by reducing the number of RF chains while maintaining high performance, making large-scale deployment of high-frequency systems more practical [201].
Nguyen et al. [118] proposed a deep unfolding framework by unrolling the iterations of Riemannian manifold optimization (Riemannian MO), specifically the MO-AltMin algorithm [202]. Instead of directly minimizing WMMSE and maximizing SE, the problem is transformed into an approximate matrix factorization problem, which is then solved using LS. This approach is computationally more efficient, as it requires fewer layers to compute the weights of the precoding matrix, significantly reducing the training time. Compared to the approach in [75], the proposed network achieves better performance while maintaining much lower time and computational complexity.
Minghui et al. [119] proposed a DL-based RSMA scheme for IRS-aided THz multi-user MIMO systems, focusing on enhancing robustness against CSI imperfections to improve system performance. The approximate WMMSE algorithm for RSMA is unfolded in this approach to boost precoding performance while reducing computational complexity. The deep unfolding algorithm processes the estimated equivalent channel through a series of layers to optimize the digital active precoding. This strategy aims to mitigate performance loss caused by model mismatches in real-world systems. By deep unfolding the approximate WMMSE algorithm, the method achieves improved precoding performance and lower computational complexity compared to the original WMMSE approach. This iterative deep unfolding process refines the precoding strategy based on the estimated equivalent channel, optimizing performance in IRS-aided THz massive MIMO-OFDM systems.
In [120], Adam et al. introduced a method to jointly optimize active beamforming, passive beamforming, and UAV trajectory to minimize power consumption while accounting for UAV jitters and imperfect hardware constraints. They designed a hybrid semi-unfolding DNN to address these constraints during the real-time CSI gain estimation for IRS and UAV links. The approach is termed “semi-unfolded” because it combines two types of networks: one based on deep unfolding and the other based on a multi-kernel convolutional LSTM (MK-CLSTM) NN. For active beamforming, the authors employed semidefinite programming to obtain a closed-form solution, which was then transformed into NN layers to replicate the iterative beamforming solution in real time. The Lagrangian functions with well-defined Karush-Kuhn-Tucker (KKT) conditions were derived to design the layers of the deep unfolded architecture. The UAV trajectory and beamforming vectors are the parameters updated within these layers. For passive beamforming, a MK-CLSTM-based NN was used. Training was done using mini-batch SGD. The authors claim that their model-driven approach achieved $99.24\%$ accuracy and is $26.21\%$ faster than the classical iterative methods, offering significant improvements in both efficiency and speed for UAV-based communication systems.
VII Sensing and Communication
The independent operation of sensing and communication as separate systems has been widely studied [203]. However, with the deployment of wireless communication services in higher frequency bands, such as mmWave and THz, spectrum congestion has become a major issue. In the 6G era, independent sensing systems experience interference from communication signals, leading researchers to recognize the shared system architecture and signal processing techniques between sensing and communication [204], [205], [206]. This realization has driven the development of ISAC, where both functionalities are merged into a unified system.
ISAC can be categorized based on sensing methods into active sensing and passive sensing [207].
  • Active sensing (Mono-static sensing): The transmitter and receiver are co-located, allowing the system to extract target parameters using both the transmitted signal and received echoes. However, this approach suffers from significant self-interference at the ISAC receiver due to the strong reflections from the transmitted signals.
  • Passive sensing (Bi-static sensing): The sensing echoes are received at a different location than the transmitter, eliminating the need for in-band full-duplex operation at the receiver. This approach is more suitable for applications where reducing self-interference is a priority.
ISAC provides a promising solution for optimizing spectral efficiency and enhancing system performance in future wireless networks. However, understanding the fundamental performance limits of ISAC systems is crucial for their practical implementation. The estimation accuracy of key sensing parameters, such as direction of arrival, Doppler frequency, velocity, and target position, was thoroughly investigated in [208]. This study provided a detailed analysis of both theoretical and practical limits of ISAC systems. The authors emphasized that the performance limits of ISAC channels cannot be derived by simply merging the existing bounds from independent sensing and communication systems. Instead, a unified framework is necessary to characterize these limits effectively. To address this, the authors proposed deriving unified Cramér-Rao bounds for ISAC, which establish the fundamental lower bounds on sensing accuracy given a specific allocation of orthogonal resources between communication and sensing. These unified expressions provide insights into how ISAC systems can optimize their resource allocation strategies to achieve superior performance in joint sensing and communication tasks.
VII-A MIMO ISAC
The integration of deep unfolding in ISAC systems has shown promising advancements in optimizing waveform design, beamforming, and communication-sensing tradeoffs. Ziwei et al. [82] applied deep unfolding techniques to optimize waveforms in MIMO radar systems, specifically focusing on minimizing the range-integrated sidelobe level. This is crucial for reducing interference and improving target detection performance. Their deep unfolding network efficiently optimizes radar waveforms, enhancing overall radar performance. In [83], the challenge of designing constant-modulus waveforms for multiuser MIMO ISAC systems was explored. Since constant-modulus waveforms help avoid signal distortion in nonlinear power amplifiers, their design is crucial. However, the non-convex nature of the constant-modulus constraint makes optimization challenging. To address this, the authors employed a deep unfolded network with a sparsely connected structure, trained in an unsupervised manner, ensuring efficient waveform design under these constraints. Hybrid beamforming in ISAC systems poses additional complexity due to the need to balance communication rates and sensing accuracy. The authors in [84] tackled this challenge by unfolding the iterative PGA algorithm. Their deep unfolding-based method accelerated convergence and enhanced overall system performance, achieving an optimal balance between communication and sensing objectives. Further, joint optimization of radar-communication beamforming was addressed in [209]. The authors incorporated symbol-level precoding with constructive interference constraints to improve energy efficiency by exploiting multi-user interference. They developed iterative algorithms to leverage problem separability, enabling parallel and recursive computation. Deep unfolding was utilized to reduce computational complexity, improve scalability, and support real-time implementation.
VII-B mmWave MIMO ISAC
mmWave MIMO is a key technology in 6G wireless communication, offering high transmission capacity and sub-meter user localization. Its large-scale antenna arrays enable precise angular resolution. Traditional localization methods estimate position-related parameters before reconstructing the user’s location. To improve this, Fan et al. proposed a fast direct localization method in [89]. They use an unfolded ADMM solver with low complexity for direct localization. The trainable parameters include $\rho$ (penalty parameter), $w$ (weight coefficient for channel gain between the base station and user), and $\tau_{1}$, $\tau_{2}$ (two positive proximal parameters).
VII-C Modern Multicarrier Modulation for ISAC
5G and beyond 5G wireless communication systems operate at carrier frequencies in the microwave regime, where radar applications are dominant. This has led to the development of joint waveforms for sensing and communication to address limited spectral resources [210]. OFDM-based radars in LTE and 5G NR mobile networks’ BS are studied for radar and sensing in [211], [212], [213]. OTFS modulation for joint radar parameter estimation and communication is explored in [214], [215], [216].
An ISAC signal processing problem was proposed in [102] to process pilot and data signals together. The optimization problem requires the simultaneous recovery of CSI, transmitted data symbols, and target information, making it non-convex. To solve this, the authors used an iterative optimization algorithm that alternates between passive sensing and communication detection. Jingcai et al. proposed a deep unfolded network for target parameter estimation in MIMO-OFDM-based ISAC systems [103]. Their method leverages the sparsity of target parameters, using a compressed sensing approach. The problem is reformulated as a least absolute shrinkage and selection operator (LASSO) problem. The fast ISTA algorithm is then unfolded, with learnable parameters enabling high-precision solutions with low computational complexity.
The problem of estimating parameters of moving targets in high-speed scenarios using OTFS modulation for ISAC systems is explored in [104]. The target parameter estimation is reformulated as a compressed sensing problem. For the deep unfolded network, ADMM iterations are unfolded into the network’s layers.
VII-D IRS-Assisted ISAC
Liu et al. explored beamforming design for an IRS-assisted ISAC system with bounded angle uncertainty of sensing targets and statistical CSI uncertainty related to user equipment in [117]. They formulated a bi-level optimization problem, unfolding the projection method and Riemannian MO approach. These methods handle the maximum transmit power and unit-modulus constraints. The bi-level problem consists of two parts: the lower-level problems focus on identifying the worst-case CSI and target sensing angles, while the upper-level problem addresses beamforming design and optimization of Lagrange multipliers based on the given CSI and target angles. A double-loop deep unfolding NN is applied, unfolding the iterative gradient descent into a multi-layer structure, with each layer containing an upper-level and a lower-level loop. Trainable step sizes and offset parameters are introduced as network parameters.
VIII Decoding in Error Correcting Codes
Error control codes are essential for reliable and efficient communication. During data transmission, bits may flip due to channel noise, causing errors. Error correction coding helps detect and correct these errors at the receiver. Claude Shannon revolutionized channel coding with his landmark paper [217].
Mobile communication has evolved through several generations with significant technological advancements. In the first generation (1G), communication was primarily analog, leaving no need for channel coding. With the second generation (2G), convolutional and block codes were introduced to compress voice signals and convert them into digital form. Turbo codes were adopted as standard coding schemes in the third (3G) and fourth (4G) generations, offering near-Shannon limit performance with moderate decoding complexity [218]. Low-density parity-check (LDPC) codes, first introduced in the 1960s, were later rediscovered and extensively studied [219], [220], [221]. They are widely used in 4G, 5G, and beyond. LDPC codes are represented using a bipartite graph, placing them in the “codes on graphs” category, designed for capacity-approaching codes with iterative decoding algorithms. Since LDPC codes use sparse graphs, the number of edges grows linearly with block length, making them low-density and less complex [222]. Polar codes, introduced by Erdal Arıkan in 2009, are a modern class of error-correcting codes designed to achieve the capacity of symmetric binary-input discrete memoryless channels [223]. They are based on channel polarization and are known for their capacity-achieving property, meaning they can approach the Shannon limit as block length increases.
BP is widely used to decode modern error correction codes [224]. This decoder is constructed from the Tanner graph, introduced by Michael Tanner in his landmark paper [225]. He developed the min-sum and max-sum algorithms, which are derived from the generic sum-product decoding [226]. Polar codes use successive cancellation decoding, where received bits are decoded sequentially by leveraging frozen bits and previously decoded bits. More advanced techniques, such as successive list decoding assisted by cyclic redundancy check (CRC), further improve performance [227].
In practical wireless communication, channels often deviate from mathematical models or are not precisely known at the decoder. In such cases, a DL-based decoder offers a robust and adaptive solution. Additionally, model-driven DL is well-suited for providing an interpretable analysis of iterative decoding algorithms. Wei et al. [228] proposed an unfolded ADMM-based penalized algorithm for binary linear codes. Their approach introduces two networks that transform the ADMM penalty parameters $\{\mu,\alpha\}$ into iteration-dependent parameters, along with an adjustable penalty function. Each stage of the learnable ADMM decoding network (LADN) corresponds to one ADMM iteration, where iteration-dependent parameters enhance convergence and reduce the number of iterations. The key idea is that increasing the number of learnable parameters improves generalization capability. For the second network, a piecewise linear penalty function with adjustable slopes is designed to suppress pseudo codewords effectively, improving decoding performance beyond conventional $\ell_{1}$ and $\ell_{2}$ norm functions. Parameter updates are trained using backpropagation via the Adam optimizer. A customized loss function based on MSE ensures effectiveness at every iteration by considering outputs from all layers. The authors claim that their decoder outperforms the conventional ADMM decoder. Additionally, decoding irregular binary LDPC codes via unfolded ADMM is explored in [229].
In [48], the authors propose the TurboNet architecture, which combines DL with the max-log MAP algorithm. While most existing model-driven decoding algorithms rely on BP and are iterative, the authors enhance the traditional MAP algorithm by unfolding it, with each iteration represented as a DNN decoding unit. As a result, the soft-input soft-output decoder, which traditionally uses the max-log MAP algorithm and extrinsic LLRs, is replaced by two subnets based on a neural log-MAP algorithm for each iteration.
In [230], the authors tackle the decoding challenges of marker codes in insertion and deletion channels, with a focus on overcoming CSI uncertainty. To improve detection performance, they propose DL techniques, implementing both RNNs and a deep unfolded network. The forward-backward algorithm is unfolded into multiple layers within the network to enhance the overall decoding process.
Gao et al. [231] have unfolded the factor graph of the BP decoder using a one-bit quantizer for polar codes. In [232], a double-loop iterative decoding algorithm for LDPC codes is unfolded within a penalty dual decomposition framework. Additionally, by unfolding the iterative decoding process between check nodes and variable nodes into a feed-forward propagation network, a deep unfolded normalized min-sum LDPC decoder is derived in [233].
Tanner-graph-based transmission schemes for massive multiple access, incorporating NOMA and grant-based random access, are discussed in [234]. To enhance decoding efficiency, a message-passing decoder is unfolded for both transmission schemes, allowing the decoder to improve as more information about the transmitted data is learned during the training epochs. In [235], theoretical results for the unfolded BP decoder are presented. The authors define the generalization gap of a decoder as the difference between the empirical and expected bit-error rates. They provide a set of theoretical results that bound this gap and examine its dependence on decoder complexity, which is influenced by code parameters (such as block length, message length, and variable/check node degrees), the number of decoding iterations, and the size of the training dataset.
IX Power allocation
In a multi-antenna wireless communication system with multiple users and limited resources, ensuring reliable transmission of massive amounts of data presents significant challenges. One of the primary constraints is the power budget, which must be allocated efficiently among data streams to maintain a high quality of service. The water-filling algorithm is a widely recognized technique for optimal power allocation across subchannels, aiming to maximize the sum capacity or sum rate. Since power allocation directly impacts channel capacity, it is formulated as an optimization problem, where maximizing the capacity rate leads to improved system performance. The power allocation problem can thus be formulated as an optimization problem as shown in (6) below:
$\max_{\left\{P_{i}\right\}}$ $\frac{\sum_{i=1}^{K}w_{i}R_{i}(P_{i})}{\sum_{i=1}^{K}P_{i}}$ (6)
s.t. $R_{i}(P_{i})\geq R_{\min,i},\quad\forall i,$
$\sum_{i=1}^{K}P_{i}\leq P_{\text{total}},$
$P_{i}\geq 0,\quad\forall i.$
where $K$ is the number of users, $w_{i}$ is the weight associated to user $i$ decided by user priority or fairness, $R_{i}(P_{i})$ is the rate achieved by user $i$ given the power $P_{i}$ allocated to them, $R_{\text{mini},i}$ is the minimum required rate for fairness and $P_{\text{total}}$ is the sum of all users’ power values.
Several classical approaches, such as the Dinkelbach algorithm, Lagrange dual decomposition, interference pricing, successive convex approximation (SCA), WMMSE, and matching theory, have been employed for power allocation [122, 236, 237, 238, 239, 240]. However, a major drawback of these methods is their reliance on iterative algorithms, which are computationally intensive and result in long runtimes, ultimately degrading system energy efficiency. The deep unfolding approach mitigates this issue by reducing the number of iterations required to achieve a near-optimal solution, thereby enhancing computational efficiency and energy savings.
Wireless networks expand in terms of spatial coverage, but traditional DNN architectures, such as multi-layer perceptrons and CNNs, struggle to capture the inherent topology of these networks. To address this limitation, Chowdhury et al. [241] propose leveraging GNNs to parameterize the power allocation function. In their approach, the WMMSE algorithm is unfolded into layers, resulting in a distributed, efficient, and scalable solution. The unfolded WMMSE method integrates domain-specific elements with trainable components, which are parameterized by GNNs to exploit the natural graph structure of wireless communication networks. The objective is to map the channel matrix and approximate optimal power allocation by training the GNN using a block coordinate descent algorithm.
In [242], Li et al. unroll the SCA algorithm into NN layers for energy-efficient power allocation in wireless interference networks. The proposed unfolded algorithm is highly adaptable, as it generalizes across varying network sizes and channel distributions while also enabling transfer learning between different path loss models. Given that the system operates in a multi-user wireless interference network, it can be naturally modeled as a graph neural network GNN. The power allocation policy is formulated based on the CSI matrix ${\bf{H}}$ and the total power constraint $P_{\text{total}}$, within a layered architecture incorporating trainable weights.
Hu et al. [243] propose an unsupervised power allocation optimization approach for ad-hoc networks by integrating attentive graph representation with deep unfolded WMMSE. The key advantage of attentive graph representation is its ability to allow network nodes to selectively focus on the most relevant connections during learning, rather than treating all links equally. The power allocation framework is built upon this attentive graph representation, as described in [244, 245]. Meanwhile, the deep unfolded WMMSE maps the iterative steps of the WMMSE algorithm into neural network layers within an unsupervised learning paradigm. This approach effectively combines deep unfolding with graph representation learning, significantly reducing the search space for the objective function. As a result, the model achieves improved interpretability and accelerated training while eliminating the need for supervised data.
X Physical Layer Security
Physical layer security is a critical research area aimed at ensuring secure and reliable transmission between the transmitter and receiver while preventing eavesdropping and impersonation attacks. Due to the broadcast nature of wireless communication, adversarial users pose a significant threat. These adversaries generally fall into two categories: eavesdroppers (or middleman attackers), who attempt to extract information from an ongoing transmission without detection, and jammers, who deliberately degrade the signal quality to prevent the receiver from successfully decoding the intended message [246]. Developing robust security mechanisms at the physical layer is essential to mitigate these threats and enhance communication reliability.
The fundamental network model for analyzing security and confidentiality is a three-terminal system, as illustrated in Fig. 8. In this setup, the legitimate transmitter (Alice), the intended receiver (Bob), and the unauthorized eavesdropper (Eve) are equipped with $N_{\text{t}}$, $N_{\text{r}}$ and $N_{\text{e}}$ antennas, respectively. As introduced by Wyner in his seminal work on wiretap channels [247], this model captures the imperfections introduced by the communication channel. Alice transmits an information signal, $\bf{x}$, a random vector intended for Bob, over the main channel, which is modeled as a discrete memoryless channel. However, due to the broadcast nature of wireless communication, the signal received by Bob, $\bf{y}$, is also partially intercepted by Eve through an additional wiretap channel, resulting in her received signal, $\bf{z}$, as shown in Fig.  8.
Figure 8: General wiretap channel [246].
Traditionally, secret key-based encryption has been the standard approach to ensuring message confidentiality, where the encryption key is shared exclusively between the transmitter and the intended receiver. As defined in [248], a secrecy system can be abstractly described as a set of transformations that map a set of possible messages to a set of possible cryptograms. Each transformation corresponds to an encryption process that utilizes a unique, secure key to encipher the message, ensuring that unauthorized entities cannot decipher the transmitted information.
The primary objective of a secure communication system is to maximize the secrecy rate, ensuring that the transmission rate to the intended user remains high while minimizing information leakage to the eavesdropper. The secrecy capacity $C_{s}$ is defined as the maximum rate $R_{s}$ at which secure communication can be achieved, guaranteeing that the eavesdropper receives negligible information about the transmitted message. This capacity represents the fundamental limit of secure data transmission, balancing efficiency and confidentiality in the presence of adversarial interception. The secrecy capacity can be expressed as:
$C_{s}=\max_{P_{X}}\left\{I\left({\bf{X};\bf{Y}}\right)-I\left({\bf{X};\bf{Z}}\right)\right\}$
where, $I(\bf{X};\bf{Y})$ is the mutual information between the transmitter and the intended user, and $I(\bf{X};\bf{Z})$ represents the information leakage to the eavesdropper. The transmitter can maximize the secrecy capacity by carefully choosing the probability distribution $P_{X}$ for the transmitted signal $\bf{x}$ to maximize the difference $I(\bf{X};\bf{Y})$-$I(\bf{X};\bf{Z})$ i.e, the rate $R_{s}$.
Jamming attacks pose a serious threat to the reliable operation of wireless communication systems [249, 250]. A jamming attacker deliberately transmits interference or noise on the same frequencies as legitimate users. This disrupts communication channels, reduces signal quality, and can block data transmission entirely. The interference makes it harder for users to decode messages, leading to connection losses or forced retransmissions. As a result, network efficiency decreases, and latency increases. Marti et al. in [251] explored solutions to combat smart jamming in massive multi-user MIMO systems. In these systems, jamming attacks can selectively interfere at specific times to avoid detection. The authors proposed an optimization-based approach for joint jammer estimation, channel estimation, and data detection, using the BS’s multi-antenna configuration. This method does not require prior knowledge of the jammer’s behavior, making communication more resilient even during sporadic or bursty jamming. Using a deep unfolding approach, the authors developed two iterative algorithms. The first is mitigation estimation and detection (MAED), which uses forward-backward splitting for estimation. The second is soft-output MAED (SO-MAED), an enhanced version that applies deep unfolding to optimize parameters. SO-MAED improves detection accuracy, especially in high-interference environments, and does not require knowledge of the attack type. Error precision and gradient step sizes are key parameters for improving learning stability.
XI Joint Techniques
This section reviews different joint techniques that are aided by deep unfolding method.
XI-A Joint Detection and Channel Estimation
An iterative method for estimating joint channel parameters and selecting symbols using the EM algorithm is presented in [252]. This method performs symbol detection in the presence of Gaussian noise and intersymbol interference in a dispersive channel. The authors model the missing data using a Markov chain, and channel parameter estimation follows the ML criterion. The Markovian properties of the channel state sequence allow for calculating the required likelihoods using the forward-backward algorithm. In [253], the authors explore a V-BLAST-type combination of OFDM with MIMO to improve SE and multi-user downlink throughput. They propose a joint data detection and channel estimation algorithm for MIMO-OFDM that integrates the QR decomposition M-algorithm and Kalman filter. The M-algorithm, combined with QR decomposition, reduces the computational complexity of tree-based search methods [254].
As mentioned in [6], the data-driven approach uses a fully connected network, where each sub-block of the OFDM transceiver is designed. Gao et al. [255] proposed a deep unfolding approach to address the issue of DNN performance being dependent on large training datasets. In contrast to the approach in [6], where a DL architecture completely replaces the conventional modules, the ComNet architecture shown in Fig. 9 incorporates both channel estimation and detection modules. Both modules use conventional communication solutions as inputs to the network. The DL network then refines these coarse inputs. For example, the LS channel estimate, $\bf{\hat{h}_{\text{LS}}}$, is refined using pilot symbols ${\bf{x}}_{p}$ and corresponding received symbols ${\bf{y}}_{p}$ as inputs to the LS sub-network within the channel estimation module. The channel estimate, $\bf{\hat{h}_{\text{LS}}}$, is passed from the LS sub-network to the LS Refine Net sub-network to generate an accurate channel estimate, $\bf{\hat{h}}$. A single-layer DNN is used to design the LS Refine Net. The refined channel estimate, $\bf{\hat{h}}$, from the LS Refine Net is then passed as input to the ZF detector sub-network in the signal detection module. The ZF detector sub-network is also fed with the received data vector, $\bf{y_{\text{D}}}$, and the LS channel estimate, $\bf{\hat{h}_{\text{LS}}}$. This is followed by the ZF Refine Net. The refined channel estimate, $\bf{\hat{h}}$, the received data vector, $\bf{y_{\text{D}}}$, and the ZF estimate, $\bf{\hat{x}_{\text{ZF}}}$, from the ZF detector sub-network are passed to the ZF Refine Net to estimate the transmit binary data, $\hat{b}$. The short dotted line from the LS sub-network to the ZF detector sub-network in Fig. 9 provides a robust candidate for the recovered binary symbols, $\hat{b}^{{}^{\prime}}$, at deQAM. Both the channel estimation and detection modules are trained using an $\ell_{2}$ loss between the estimated binary data, $\hat{b}$, and the raw transmit binary symbol, $b$. Simulations have shown that this design outperforms both LMMSE and FC-DNN [6].
Figure 9: ComNet architecture [255].
An artificial intelligence-aided receiver for a cyclic prefix-free OFDM system with joint detection and channel estimation is explored in [256]. Zhang et al. [257] unfolded the OAMP algorithm for generalized frequency division multiplexing (GFDM) systems with low-resolution analog-to-digital converters. In [258], a general EM algorithm is derived and unfolded to address the impact of channel estimation errors on signal detection. The expectation step estimates the CSI, while the maximization step detects the transmitted signal using the estimated CSI. A modified, trainable projected gradient detector is used in the maximization step. In [259], an ADMM-based unfolded algorithm is proposed for robust MIMO detection, incorporating channel estimation errors. To reduce error propagation due to channel aging, a robust data-aided Kalman filter-based channel tracking method is developed. The pre-estimated data symbols are used to improve channel estimation accuracy, enhancing robust detection performance.
Recently, the energy harvesting technique has gained significant attention as a promising feature for green communication networks. Ambient backscatter communication is an energy-efficient technology that allows devices to transmit data by reflecting existing ambient signals, such as those from Wi-Fi routers or cellular networks, instead of generating their own radio waves. A backscatter device modulates the signal over the ambient radio frequency without an active radio frequency component. However, the backscatter link is weaker than the direct link, which limits its coverage and data rate [260]. In [261], Wang et al. explored an IRS-based ambient backscatter communication system. The proposed network consists of two parts. The first subnetwork calculates the IRS phase shift matrix for binary symbols. The phase shift matrix weights, or phase angle vectors, are updated using a DNN training tool. These updated phase angle vectors are then used to compute the received signal, which is passed to the second subnetwork, the detection module. The EM algorithm is unfolded for symbol detection.
XI-B Joint Activity Detection and Channel Estimation
In [262], Ke et al. addressed massive access in massive MIMO systems by introducing an adaptive scheme for active user detection and channel estimation using compressive sensing techniques. This scheme takes advantage of the sparse traffic patterns of many user devices and the sparse nature of massive MIMO channels in the virtual angular domain. The access latency is reduced due to the use of non-orthogonal pseudo-random pilots for uplink communication. The problem of detecting active users and estimating their channels is formulated as a generalized multiple measurement vector compressive sensing problem. A generalized multiple measurement vector AMP (GMMV-AMP) algorithm is designed to leverage the structured sparsity in the uplink channel matrix. Additionally, a Turbo-GMMV-AMP algorithm is developed to exploit channel sparsity in both the spatial and angular domains. In [263], an EP-based joint active user detection and channel estimation technique for mMTC networks is proposed. The EP algorithm iteratively minimizes the Kullback-Leibler divergence between the target and approximate distributions via moment matching, efficiently finding a tractable close approximation of the target distribution.
Grant-free random access is used for active device detection at the BS to address the latency issue in massive access. The active device detection is formulated as a sparse signal recovery problem. To improve the performance of massive device detection, Qiang et al. [43] proposed a deep unfolded network called AMP-Net, as shown in Fig. 10. The basic principle of AMP-Net is based on the AMP algorithm. The input variables passed through the $T$ layers include the state evolution parameter, $\tau^{2}$, the device state matrix $\bf{X}$, the received signal $\bf{Y}$ at the BS, and the residual matrix $\bf{R}$. The pilot vectors are denoted by $\bf{a_{\text{n}}}$. Each layer contains an MMSE estimator, $\eta_{t}(\cdot)$. The set of trainable parameters consists of four scalar values. The parameters $\lambda_{n}^{t}$ and $\gamma_{n}^{t}$ estimate the activity probability and channel variance, respectively. The parameter $\theta_{n}^{t}$ regulates the denoiser, while $\mu_{n}^{t}$ controls the state evolution parameter, helping facilitate an accurate prediction of the detected target.
Figure 10: The structure of AMP-Net [43].
Traditional orthogonal multiple access fails to achieve massive connectivity for mMTC systems. Random and sporadic data traffic means that only a small fraction of users access the BS during uplink transmission. The joint channel estimation and active user detection problem for uplink mMTC is investigated in [264]. The authors developed a pilot transmission scheme for nonorthogonal identification. Due to the sporadic activity of users, where only a small fraction are active at any time, traditional orthogonal pilots are inefficient. Instead, nonorthogonal pilots are used, and the joint estimation and detection problem is formulated as a compressive sensing problem. The work considers single and multiple antenna configurations at the base station, modeling them as single measurement vector and multiple measurement vector problems, respectively. The authors designed a deep unfolding model based on the ADMM algorithm. For ADMM, the trainable parameters include: (1) an auxiliary variable to compute matrix inversions, (2) a regularization parameter that controls the penalty term in the ADMM optimization, influencing the balance between sparsity and accuracy in the recovered signal, and (3) a threshold parameter in the shrinkage function that determines the sparsity of the output in each iteration. In [265], two deep unfolded neural network architectures are proposed for sporadic user detection and channel estimation in mMTC. The first architecture unfolds a linearized ADMM, applied to a constrained convex problem. Linearized ADMM addresses computationally intractable block coordinate sub-problems by executing a single iteration of the proximal gradient method for each subproblem instead of solving them explicitly [156]. The second architecture unfolds a vector AMP with a denoiser based on ISTA. This unfolded network reduces complexity and improves accuracy compared to the original algorithm.
In [266], a deep unfolded algorithm is proposed for the joint design of compression modules, quantization modules at the APs, and the decompression and detection modules at the CPU for cell-free massive MIMO systems with capacity-limited fronthauls. The objective is to address activity detection in an end-to-end learning framework. Ma et al. [267] investigated joint device activity and data detection in mMTC. They proposed a one-phase non-coherent scheme, where data bits are embedded in the pilot sequences. The BS simultaneously detects active devices and their embedded data bits without explicit channel estimation. The AMP algorithm is unfolded to enhance detection performance by effectively exploiting the pilot activity correlation. In [268], a contention-based asynchronous random access scheme with data length diversity is considered for mMTC. The paper unfolds AMP with backpropagation to address the composite problem of activity detection, channel estimation, and data recovery. Zou et al. [269] formulated the joint activity detection and channel estimation problem as a group-row-sparse matrix recovery problem. A proximal gradient-based unfolding NN is developed to parameterize the algorithmic iterations and solve this problem. In [270], compressive sensing recovery algorithms are used to address the joint active device detection and channel estimation problem for uplink transmission in grant-free based vehicular networks. The ADMM algorithm is unfolded to improve convergence rate and recovery accuracy. Shi et al. [271] considered grant-free massive access in an IoT network. A large number of single-antenna IoT devices sporadically transmit non-orthogonal sequences to a multi-antenna BS. An unfolded DNN framework is developed using the multidimensional shrinkage thresholding operator inherited from the ISTA structure. This network simultaneously achieves low computational complexity and high robustness for solving the joint activity detection and channel estimation problem.
Cell-free technology for massive MIMO enhances mmWave signal coverage [272]. Using deep unfolding techniques, the authors in [273] developed a grant-free access protocol based on mmWave cell-free massive MIMO. The proposed hybrid structure leverages learned vector AMP to reduce the need for prior channel information and achieve faster convergence.
XI-C Joint Channel Estimation and Precoder Design
A codebook design for channel estimation in mmWave communications with a hybrid precoding structure is studied in [274]. To evaluate the performance of a codeword, the authors propose a metric called generalized detection probability (GDP), which considers both the beam pattern’s flatness and the maximum transmission power. Yuan et al. [275] introduced a two-stage wideband hybrid beamforming scheme for THz wireless systems. The model features a multi-antenna BS communicating with a multi-antenna user over frequency-selective fading. Additionally, a probabilistic robust hybrid beamforming scheme is designed to mitigate channel estimation errors.
(a) Long-term
(b) Short-term
Figure 11: JCEHB architecture [276].
Instead of modeling and optimizing hybrid beamforming and channel estimation separately, Kang et al. [276] proposed the joint channel estimation hybrid beamforming (JCEHB) algorithm. The sum rate is maximized for TDD massive MIMO system in a mixed timescale framework. The algorithm’s architecture is shown in Fig. 11. Observe from Fig. LABEL:sub@fig:fig1, the deep-unfolding NN architecture consists of two parts : channel estimation and hybrid beamforming. For channel estimation, the pilot training module is designed by unfolding the recursive least squares (RLS) algorithm. The SSCA algorithm is unfolded to replace the high complexity computation operations in hybrid beamforming. The long-term analog beamforming matrices, are optimized offline with CSI statistics that is constant over long time slots or a superframe. After fixing the analog beamformers, the deep-unfolding NN for channel estimation is trained at each time slot to optimize the digital beamformers, as shown in Fig. LABEL:sub@fig:fig2. In contrast to long-term analog beamforming, short-term digital beamforming matrices are optimized every time slot using real-time low-dimensional effective CSI. Analog beamforming matrices are optimized offline using CSI statistics, while digital beamforming matrices are updated within each time slot based on estimated low-dimensional CSI. The authors claimed that the hybrid approach in a mixed (long-term and short term) timescale framework reduce signaling overhead as compared to separate design of NN in a single timescale.
An unfolded WMMSE-based NN for massive multi-user MIMO under imperfect CSI is studied in [277]. The proposed network enhances performance with low complexity by leveraging CSI statistics for near-optimal precoder solutions, accelerating convergence. In [278], an unfolded sparse Bayesian learning algorithm is derived to capture channel sparsity in a mmWave massive MIMO system with hybrid beamforming. Similarly, [279] proposes a learning-based channel semantic acquisition and beamforming solution for cell-free massive MIMO. A multi-layer perceptron estimates the channel, while a successive over-relaxation-based linear beamforming scheme reduces estimation errors in hybrid beamforming. Wideband precoding under imperfect CSI is discussed in [280]. A robust frequency-selective precoding algorithm maximizes the expected weighted sum rate while considering per-antenna power constraints. This algorithm is unfolded to reduce computational complexity. To improve beamspace channel estimation, [281] introduces a deep unfolding algorithm using the top $\ell_{K,2}$ norm. This norm, a variation of the traditional $\ell_{2}$ norm, focuses on the largest elements in the sparse beamspace channel vector [282]. It helps determine sparsity levels, reducing noise and computational complexity. The iterative trimmed-ridge regression algorithm is unfolded to allow parallel computation. These algorithms are applicable to both narrowband and wideband massive MIMO systems, ensuring versatility in channel estimation.
XI-D Joint Detection and Decoding
An adaptive Bayesian receiver using sequential Monte Carlo for signal detection and decoding in fading channels with known channel statistics is proposed in [283]. The transmitted signal is treated as missing data, and multiple signal samples are sequentially estimated based on observed signals. These estimated sequences approximate the Bayesian estimate of both transmitted signals and channel states. Adaptive receiver algorithms are developed for both uncoded and convolutionally coded systems. El Gamal et al. [284] present an iterative receiver for joint detection and decoding of CDMA signals. The multiuser detector follows the MMSE criterion. A modified symbol-aided demodulation technique extracts soft information from previous decoding iterations to improve channel estimates, enhancing joint detection and decoding efficiency.
Traditionally, channel estimation, signal detection, and channel decoding in a MIMO receiver are performed sequentially and independently. This disjoint approach leads to suboptimal performance. To address this, Sun et al. [285] propose receivers that integrate these functions, reducing error propagation and processing delay. Their design targets MIMO ultra-reliable low-latency communication (URLLC) systems. The model uses short channel codes and a limited number of pilot symbols for channel estimation. The authors employ the ADMM algorithm to solve large-scale non-convex optimization problems, deriving closed-form solutions for each iteration. Two NNs, JCDDNet-G for Gaussian channels and JCDDNet-S for sparse mmWave channels, are built by unfolding the ADMM algorithms as shown in Fig. 12. Simulations show that the proposed trainable JCDD receivers significantly outperform traditional turbo receivers.
Figure 12: JCDDNet-S/G architecture [285].
In [286], Zhang et al. propose a receiver framework based on a deep unfolded NN structure. The joint signal detection and channel decoding modules are modeled as multi-layer deep feed-forward networks. Signal detection is achieved by unfolding the EP algorithm, while channel decoding is based on unfolded turbo decoding. Building on this work, Zhang et al. develop an unfolding-based framework for MIMO detectors in [287]. This framework adaptively determines trainable parameters, allowing the detector to adjust effectively to varying channel conditions.
XII Challenges in Developing Deep Unfolding Models and Future Research Scopes
The extensive review of recently emerging techniques and applications of deep unfolding have demonstrated its potential in addressing complex signal processing problems with reduced complexity and runtime. Despite these advantages, the adoption of deep unfolding in practical scenarios remains limited. In this section, we outline the main challenges in developing deep unfolding models, explore the reasons behind these challenges, and propose potential improvements to make deep unfolding a more powerful and widely applicable technique in wireless communications.
XII-A Limitations of Deep Unfolding Techniques
Although deep unfolding has found applications in various design problems, it has limitations. Most of these limitations arise from the fact that deep unfolding is not an independent design methodology; it must be built upon existing conventional optimization frameworks. Consequently, developing deep unfolding solutions faces the following challenges:
(C1) Contrasts between conventional optimization and deep unfolding: There is a notable contradiction between the development methodologies of conventional optimizers and deep unfolding models. Most conventional solutions for transceiver design and signal processing in wireless communications are based on convex optimization techniques. However, many wireless design problems are inherently non-convex and highly coupled due to factors such as fractional SINR, constant-modulus constraints (e.g., for analog beamforming), and mixed-integer variables (e.g., in symbol detection or antenna selection). These challenges are typically addressed using a two-layer approach. In the first layer, design variables are decoupled through alternating optimization (AO) or block coordinate ascent/descent, breaking the original problem into smaller sub-problems, each focusing on a subset of variables that are solved alternately. In the second layer, techniques such as SCA or the ADMM are applied to solve each sub-problem, often introducing additional variables. Although this two-layer approach transforms the original non-convex problem into a series of convex ones that can be solved using convex optimization tools like CVX, it significantly increases the number of variables and sub-problems and leads to nested loops in the solution process. These loops, corresponding to the two layers, must iterate until convergence, adding considerable complexity. In contrast, deep unfolding models typically approach the problem as a unified whole, solving all variables in parallel, facilitating the development of unfolding architectures. This fundamental difference makes it challenging to develop suitable algorithms for unfolding. Additionally, the increased complexity in variables and sub-problems inherent in conventional methods further complicates the design of deep unfolding frameworks, directly contributing to their overall complexity.
(C2) Challenges in unfolding principle algorithms: While deep unfolding typically enhances conventional iterative algorithms, not all can be effectively unfolded. This limitation arises from the core principle of deep unfolding, which structures the iterations of an iterative algorithm as layers within a DNN. Complex operations like matrix inversion, matrix decomposition, or the use of convex optimization tools such as CVX or YALMIP (with solvers like MOSEK, SeDuMi, etc.) are difficult to implement in a DNN framework. In large-scale systems such as massive MIMO, even matrix multiplications result in high complexity for unfolding models.
Another challenge is parameter generalization, particularly with respect to the number of layers in the unfolding model. In iterative algorithms with nested loops (discussed in C1), both inner and outer loops run until convergence, but deep unfolding models require the number of layers to be predetermined before deployment. This is typically done through empirical tuning, which may not provide optimal values for all channel realizations. As a result, convergence is not guaranteed across different scenarios, potentially leading to performance degradation. Proper tuning is essential to achieve a favorable performance-complexity tradeoff in deep unfolding models.
(C3) Challenge in handling design constraints and objectives: One of the biggest challenges in developing deep unfolding models lies in managing the design constraints formulated in the problem. These constraints require the DNN layers’ outputs to meet specific requirements, such as total power, QoS, or other element-wise constraints. Simple constraints, like power or constant modulus, can often be enforced through normalization [118]. For instance, the $tanh$ activation function can be employed to impose constant-modulus constraints on analog beamforming coefficients or sensing waveforms, similar to black-box DL architectures. However, QoS constraints are significantly more complex. While they resemble activation functions, none of the existing activation functions can adequately manage such intricate requirements. One approach is to incorporate these constraints into the design objective, but this method often results in a problem that is not equivalent to the original constrained one, making the solution potentially infeasible. Properly modeling such constraints within a deep unfolding DNN remains highly challenging, hindering the development of more advanced deep unfolding models.
Not only do complex constraints pose challenges for developing deep unfolding models, but intricate objective functions also present difficulties. Generally, deep unfolding models are trained based on the objective function. For instance, unfolding architectures like DetNet [49], FSNet [152], and ScNet [288] for symbol detection are trained to minimize the Euclidean distance between the model’s output (symbol vector) and the true transmitted symbol vector. When labels are unavailable, as in beamforming problems, unfolding networks use the objective function as the loss function for minimization or its negative for maximization problems. Typical deep unfolding techniques [84, 75] require calculating the gradient of the objective function. However, this becomes challenging when the objective function is a complex, high-order, or fractional function of the design variables, especially when those variables are vectors or matrices. Therefore, more dynamic methodologies are needed for developing unfolding models that can handle such intricate functions efficiently.
(C4) Need for black-box models in sophisticated mappings: While deep unfolding is powerful, certain inference tasks still require black-box DNN models, especially when there is no clear relationship between the input data, output solution, and performance metrics. For instance, in multi-modal sensing-assisted beam tracking, tasks may involve complex mappings from environmental data—such as sensory information from sensors and cameras—to the optimal beam [289, 290]. In such cases, there is no explicit relationship between environmental images and the optimal beam, and standard communication metrics like throughput or SINR cannot be directly expressed as functions of these images. These problems cannot be formulated as optimization tasks solvable with principle algorithms, making it difficult to develop unfolding models that map environmental information to beam patterns or blockage states.
The above three challenges C1C4 also outline the scenarios wherein deep unfolding models cannot be built or does not offer much gain in performance-complexity tradeoff. These include when there is lack of suitable conventional iterative algorithms (C1, C2, C4) or when the design problems are highly complicated with constraints and objective function (C3). It can be seen that deep unfolding is not a game changer that will completely replace conventional principle algorithms or black-box ML models. However, the applicability of deep unfolding can be enhanced by overcoming those challenges, as will be discussed next.
XII-B Development of Deep Unfolding Techniques
XII-B1 Coexistence of Deep Unfolding Models, Conventional Solvers, and Black-Box Models
Deep unfolding doesn’t need to fully replace conventional algorithms or black-box ML models. Instead, these approaches can coexist, leveraging their combined strengths. In hybrid designs, deep unfolding can (i) reduce the number of sub-problems, (ii) lower the complexity of solving each sub-problem, and (iii) enhance the learning process in black-box models. Benefits (i) and (ii) correspond to challenges C1 and C2, while (iii) relates to C4.
Benefit (i) arises because deep unfolding can rely on techniques like PGA/PGD, which avoids transforming the original problem into convex sub-problems. Benefit (ii) comes from the fact that unfolded models typically exhibit lower complexity compared to methods like SCA, and their runtime is usually shorter than that of solvers like CVX. For example, in the AO framework (discussed in C1), the most challenging sub-problem often dominates the overall complexity. By applying an efficient deep unfolding model to this sub-problem, the overall efficiency improves, while simpler sub-problems can be handled by conventional solvers like CVX. Benefit (iii) is frequently seen when deep unfolding is used to preprocess input data for black-box models, as in multimodal sensing applications. The combination of deep unfolding structures with conventional solvers and black-box models leads to simpler yet more efficient designs.
XII-B2 Deep Unfolding for Multi-Objective Learning
Most existing deep unfolding models are developed for single-objective designs, such as maximizing throughput, SINR, or minimizing BER. However, in many scenarios, system designs involve multiple objectives. Traditional approaches often formulate these as constrained problems, such as maximizing energy efficiency under a QoS constraint (e.g., throughput or SINR). In ISAC designs, weighted sums of communication and sensing performance metrics are used when the goal is to balance both functions without prioritizing one over the other. The complexity grows in multi-static ISAC systems, where distributed sensing receivers have different performance metrics depending on their functions, such as SCNR, mutual information, Cramér-Rao lower bound, or probability of detection.
These designs are often difficult to solve using traditional constrained formulations, as the problem can easily become infeasible if constraints are not properly tuned. However, by incorporating all design metrics into the objective function with appropriate weights, to form a multi-objective design, the feasibility of the problem can be significantly improved [291, 84]. This multiobjective formulation also makes it easier to develop unfolding models, optimizing multiple objectives simultaneously, and improving performance trade-offs. However, developing unfolding models for multi-objective designs presents several challenges. First, as the number of terms in the objective function increases, training the unfolding DNN becomes more difficult. This is because the loss function is typically set to mirror the objective function (or its negative in maximization problems). Furthermore, balancing the changes of these objectives with respect to the design variables across layers is crucial [84, 292], necessitating careful design of the unfolding models.
XII-B3 Distributed Deep Unfolding Architectures
While deep unfolding models provide advantages in terms of complexity due to faster convergence compared to conventional iterative algorithms, they can still face high complexity issues due to challenge C2. This limits their applicability on platforms with constrained computational resources, such as edge devices or distributed access points (APs). Future wireless networks will advance beyond simple communications, evolving into large-scale, distributed systems that integrate advanced sensing, communications, and computing capabilities. However, most existing AI models, both data-driven and model-driven, were designed for centralized scenarios, where they are deployed at a central node with substantial computing power. These models are not well-suited for distributed scenarios, highlighting the need for the development of appropriate distributed unfolding architectures.
One potential solution is to develop distributed deep unfolding architectures. In this approach, rather than executing all the unfolded layers at every distributed node or edge device, the layers can be distributed across multiple platforms for parallel execution, thereby reducing latency. For example, to design $L$ beamformers for $L$ APs for their downlink transmission in a distributed system, a CPU with high computing power and hardware resources can handle the majority of the layers, generating high-quality beamformers. These beamformers are then sent to the APs, serving as initial solutions and inputs for smaller sub-unfolding models at the distributed APs. This ensures that reliable beamformers can be generated at the APs with only a few layers, minimizing complexity and runtime to meet the computational constraints of the APs. Conversely, in the uplink, a distributed unfolding detection model can be deployed at the APs to generate initial detected symbols, which are then forwarded to the CPU for further processing and final decision-making by the large unfolding detector. By such an unfolding framework, highly reliable symbols are generated with low complexity at distributed APs.
XII-B4 Deep Unfolding to Enhance Black-box Solution
As its name suggests, a black-box ML model operates with a black-box nature, making it unexplainable and resistant to re-structuring beyond tuning parameters like the number of layers or nodes. The performance and efficiency of such models are heavily dependent on the quality of input data. Deep unfolding can contribute by performing pre-processing steps to improve the input data quality. One example is in multi-modal sensing-assisted communication tasks, where data from distributed cameras, LiDAR, and radar sensors are used to enhance communication performance, such as in beam tracking and blockage prediction. However, significant challenges arise in collecting, transferring, and processing large sensory data. Specifically, high-resolution camera images generate considerable traffic overhead and latency, impacting real-time communication processing.
Deep unfolding models can be particularly useful in pre-processing such images. In image processing, deep unfolding has been successfully applied to tasks such as denoising, super-resolution, and image reconstruction, allowing for high-quality outcomes with fewer iterations than conventional algorithms. For example, unfolded algorithms based on the ISTA for sparse coding have been shown to significantly accelerate image reconstruction while maintaining high-quality results [293, 294, 295]. This makes deep unfolding a valuable tool for handling large-scale sensory data in real-time communication systems.
XIII Conclusions
The integration of traditional algorithmic principles with DL techniques has led to the development of deep unfolding, a method that overcomes the limitations associated with both conventional algorithms and standalone DL models in wireless communication. This article offers a comprehensive survey of deep unfolding techniques, examining their application across various facets of the physical layer in wireless communication systems. Key applications, such as signal detection, channel estimation, precoder design, sensing, communication decoding for error-correcting codes, power allocation, and security, are explored, with an emphasis on technologies anticipated for beyond-5G networks. The survey highlights the strengths of deep unfolding in addressing complex challenges, particularly in achieving low-latency, high-accuracy performance in dynamic environments. We conclude by discussing the key challenges in developing deep unfolding models, including issues related to scalability, computational efficiency, and robustness across diverse scenarios. Finally, we propose potential advancements to enhance the adaptability and resilience of deep unfolding, with the goal of solidifying its role as a robust and versatile technique for next-generation wireless communications.